### v1

# ECM: A Unified Electronic Circuit Model for Explaining the Emergence of In-Context Learning and Chain-of-Thought in Large Language Model

Abstract
Recent advancements in large language models (LLMs) have led to significant successes across various applications, where the most noticeable is to a series of emerging capabilities, particularly in the areas of In-Context Learning (ICL) and Chain-of-Thought (CoT). To better understand and control model performance, many studies have begun investigating the underlying causes of these phenomena and their impact on task outcomes. However, existing explanatory frameworks predominantly focus on isolating and explaining ICL and CoT independently, leading to an incomplete understanding of their combined influence on model performance. To address this gap, we propose the Electronic Circuit Model (ECM), which provides a foundation for developing scalable, learnable policies and improving the management of AI-generated content. Specifically, ECM conceptualizes model behavior as an electronic circuit: ICL is represented as semantic magnetic field to providing an additional voltage following Faraday’s Law, while CoT is modeled as series resistors to constrain the model output performance following Ohm’s Law. Experimental results demonstrate that the ECM effectively predicts and explains LLM performance across a variety of prompting strategies. Furthermore, we apply ECM to advanced reasoning strategy optimization on a series of tasks, such as the International Olympiad in Informatics (IOI) and the International Mathematical Olympiad (IMO), achieving competitive performance that surpasses nearly 80% of top human competitors.
Key Words: Electronic Circuit Model, Large Language Model, In-Context Learning, Chain-of-Thought
1 Introduction
Large Language Models (LLMs) have made significant strides across various tasks, including code generation, knowledge-based question answering, and complex reasoning [46, 3, 29, 32]. Unlike earlier models that showed incremental improvements, LLMs exhibit emergent capabilities that fundamentally reshape problem-solving. At the core of these capabilities are two critical mechanisms: In-Context Learning (ICL) [2, 11, 31] and chain-of-Thought (CoT) reasoning [26, 41, 5], enabling LLMs to adapt dynamically to new challenges and achieve human-comparable learning and reasoning [2, 27, 1]. Specifically, ICL triggers LLMs to generalize information based on task-specific context provided in the input, without requiring parameter updates. In parallel, CoT reasoning allows LLMs to approach complex problems stepwise, enhancing both performance and interpretability. Together, these mechanisms empower LLMs to serve as versatile problem solvers in critical fields such as healthcare and education [14, 25, 16, 18].
Despite their transformative potential, the mechanism of LLMs’ emergent capabilities remains poorly understood. Current research on ICL has uncovered preliminary evidence of how models encode and process contextual information, with studies suggesting latent representations that approximate forms of task-specific meta-learning [38, 36] or topic models [39]. Similarly, CoT research has begun to map the reasoning pathways LLMs employ, providing insights into their demonstration alignment with human cognitive heuristics [24, 37]. Others seek to interpret emergent behaviors using macroscopic mathematical models based on experimental analysis [13, 5]. In practice, ICL and CoT are often integrated, allowing models to adapt to context and solve problems step-by-step [17, 15, 30, 47, 6]. To put it figuratively, combining ICL and CoT is like planting a tree: ICL provides the rich “soil” in which the model can absorb and integrate contextual knowledge, while CoT guides its “branches” to grow step by step in a clear, structured direction. Together, they create a natural, human-like reasoning process that’s both flexible and transparent, enabling models to tackle complex tasks with greater depth and precision. Unfortunately, in the current literature, a unified understanding of the combined effect of ICL and CoT remains limited, which seriously affects correct performance interpretation and controllable result prediction and further constrains the development of principled methodologies for LLM design and deployment.
Figure 1: Glossary of Terms and Schematic of Electronic Circuit Model (ECM).
In-Context Learning (ICL) is conceptualized as a (Semantic) Magnetic Field. Demonstrations in ICL effectively provide an additional voltage by applying a magnetic field change aligned with the input query direction. Chain-of-Thought (CoT) process is analogized as series resistors in circuit, where the reasoning difficulty in CoT are divided into a sequence of sub-difficulties, analogous to resistors in a circuit that collectively increase overall resistor. Building on this, we introduce the concept of model inherent capabilities, represented as the power supply of the circuit to provide basic voltage. Model performance can then be quantified as the circuit’s output power, enabling more precise predictions of empirical accuracy. In addition, in this article, “strength” denotes the value of semantic magnetic fields, “resistance” denotes the value of resistor, and “voltage” denotes the value of power supply.
To bridge this gap, we introduce the Electronic Circuit model (ECM), a unified theoretical paradigm that reimagines model performance through the lens of physical circuit systems. Specifically, as shown in Figure 1, we try to conceptualize ICL capabilities as semantic magnetic fields, driving additional voltage gains that correspond to enhanced information absorption, while CoT challenges are modeled as series resistors, introducing predictable constraints akin to circuit losses. Based on these hypotheses, we discover Faraday’s Law of ICL and Ohm’s Law of CoT in LLMs’ reasoning process. By rigorously aligning theoretical circuit power predictions with empirical accuracy results across 3 prompting strategies, 4 task categories, and 13 different LLMs. The ECM successfully establishes a unified quantitative paradigm for predicting and optimizing the behavior of LLMs.
Moreover, based on the introduction of ECM, we further explain over 10 existing prompt optimization strategies from the perspective of optimizing the magnetic field and optimizing the resistors. And given these perspectives, some phenomena that have not been successfully explained by previous works and more excellent prompt strategies are discovered based on ECM. Notably, our results show that combined with these approaches has achieved breakthroughs in complex domains: LLMs guided by ECM-optimized strategies have outperformed over 80% of human competitors in the International Olympiad in Informatics and the International Mathematical Olympiad. Furthermore, in subjective and exploratory contexts such as academic research development, ECM-optimized strategies have driven significant gains, achieving a minimum 10% improvement in overall human scores.
In summary, our contributions are as follows:
  • To the best of our knowledge, we are taking the first meaningful step to introduce the Electronic Circuit Model (ECM), effectively offering a unified theoretical paradigm to reformulate model performance through the lens of circuit systems.
  • In this framework, we discover Faraday’s Law of ICL and Ohm’s Law of CoT in LLMs’ reasoning, and further utilize ECM to effectively predict the empirical accuracy and explain the effectiveness of 10 existing strategies from the view of optimizing the magnetic field and optimizing the resistor.
  • We utilize ECM to explain some open phenomena and to propose more excellent prompt strategies based on ECM. Combined with these approaches, we have achieved breakthroughs in mathematical and code competition and academic research development scenarios. Further, we empirically demonstrate that applying ECM-driven optimizations to LLMs leads to substantial performance gains.
2 Theoretical Model
This theoretical model shows that model performance ($P_{\text{out}}$) relies not only on basic voltage from model inherent capabilities ($\mathcal{E}_{\text{model}}$) but also on external factors like reasoning difficulty ($R_{CoT}$) and additional voltage from extra demonstrations in ICL ($\mathcal{E}_{ICL}$), which are described as follows in detail:
2.1 Model Inherent Capabilities as Power Supply
In our model, we draw an analogy by conceptualizing LLM’s inherent capabilities as a power supply that drives computational processes and reasoning tasks (Figure 1). The voltage value $\mathcal{E}_{\text{model}}$ represents the strength of the inherent capability, governing the LLM’s capacity to execute a range of tasks. Just as the voltage in an electrical circuit determines the power available to drive various components, the internal capabilities govern its ability to manage increasingly complex inputs.
2.2 Faraday’s Law of In-Context Learning
Inspired by Wang et al. [38], the process of ICL can be conceptualized as the interaction of LLM with contextual data, where it absorbs, retains, and eventually releases semantic information flow. This dynamic flow of information resembles the changing behavior of a magnetic field as its initial strength $\Phi^{B}_{0}$ decreases to zero. Building on this analogy, we propose Faraday’s Law for the ICL paradigm, where the rate of change in the “magnetic field" further induces an extra voltage:
$\mathcal{E}_{\text{ICL}}=-\frac{d\Phi^{B}(t)}{dt}=\lambda\Phi_{0}^{B},$ (1)
where $\lambda$ represents the uniform rate of decrease in magnetic field intensity. This perspective enables a deeper quantitative understanding of ICL through the concept of a “semantic magnetic field," as illustrated in Figure 1.
2.3 Ohm’s Law of Chain-of-Thought
In a parallel vein, Chen et al. [5] have discovered a combination law for LLM reasoning, which aligns with the series formula of electrical resistors after transformation (See in Appendix A). Inspired by this, we conceptualize the difficulty of CoT reasoning as analogous to series circuits. In this framework, each reasoning process introduces a distinct resistor ($R_{\text{CoT}}=\sum_{i}R_{i}$), compounding the task’s overall difficulty ($R_{i}$). As illustrated in the right panel of Figure 1, Ohm’s law for CoT can be expressed as:
$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}}{R_{\text{CoT}}+R_{0}},$ (2)
where $R_{0}$ represents the resistance of an output resistor with static resistance value in the circuit, which reflects the difficulty of reasoning conclusion. This formulation quantifies the cumulative effect of reasoning resistors, providing a modular quantitative perspective on cognitive task complexity.
2.4 Model Performance as Output Power of Bulb
Extending the conceptual model, we equate model performance to the power output of a light bulb in an electronic circuit. Task execution efficiency depends on a nonlinear interaction of internal and external factors, rather than solely on the model’s intrinsic capabilities. The output power, $P_{\text{out}}$, is expressed as:
$P_{\text{out}}=I^{2}_{model}R_{0}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{\text{CoT}}+R_{0})^{2}}.$ (3)
Here, output power represents the model’s effective performance, with higher power indicating better task execution, akin to a brighter bulb. The equation highlights that maximizing performance requires enhancing the basic voltage of power supply ($\mathcal{E}_{\text{model}}$) of LLM, minimizing reasoning resistor ($R_{\text{CoT}}$) in CoT, and optimizing the extra voltage ($\mathcal{E}_{\text{ICL}}$) provided by ICL. This synergy explains and predicts the model’s performance on complex tasks quantitatively.
Figure 2: The Experimental Results of Theoretical Model Verification on ECM.
a,
An analysis of ECM incorporating the CoT approach reveals a scatter plot illustrating the relationship between model accuracy and theoretical power with 0.8843 Spearman coefficient. b. Six similarity metrics are employed to evaluate whether context demonstrations can be interpreted as a semantic magnetic field. The results indicate that the projection length achieving the highest coefficient of 0.9106 between practical accuracy and theoretical power. c, The integrated ECM framework combines ICL and CoT. Across three retrieval strategies for ICL demonstrations, a consistent linear correlation is observed. d, Experiments on various LLMs under different power supply voltages, representing their capacities, reveal a strong linear correlation between theoretical power estimates and observed accuracy. e, Broader analyses across diverse tasks, including multi-hop knowledge reasoning, medical probing, and robotic planning, demonstrated robustness.
3 Theoretical Model Verification
3.1 CoT Meets Ohm’s Law
First, we verify whether CoT adheres to Ohm’s law. Since the electric current $I_{model}$ in Eq. 2 is not directly observable, we instead examine whether $P_{out}$ is linearly related to true accuracy, thereby indirectly validating CoT Ohm’s law. Specifically, to control variables, we remove the ICL and let the model perform zero-shot reasoning step-by-step following the instructions, which satisfies: $P_{\text{out}}=\frac{\mathcal{E}_{\text{model}}^{2}R_{0}}{(R_{\text{CoT}}+R_{0})^{2}}.$ The strong Spearman correlation coefficient between theoretical power ($P_{\text{out}}$) and observed accuracy (Acc), shown in Figure 2a, is 0.8843, with a $p$-value below 0.01. This strong correlation suggests difficulty of CoT reasoning actually exhibits resistor-like behavior, making it a reliable predictor of model performance.
3.2 ICL Meets Faraday’s Law
We conceptualize contextual demonstrations in ICL as a semantic magnetic field governed by Faraday’s Law during reasoning. Under this framework, semantic influence and proximity are quantified by projection length, as illustrated in the lower right of Figure 2b. To validate this hypothesis, we evaluate various metrics for contextual proximity between demonstrations and user requests. Our analysis (Figure 2b) shows that the Spearman correlation between the calculated power $P_{\text{model}}$ in Eq. 3 and observed spearman coefficient and R2 merely for projection length exceeds 0.85. This indicates that the geometric alignment of the demonstrations with the task request is a strong predictor of model performance, which further prove the existance of Faraday’s Law of ICL.
3.3 ICL and CoT Satisfy a Unified ECM
In an unprecedented integration, we unify ICL and CoT within a single ECM, offering a comprehensive framework to explain their interaction and collective impact on model performance. We rigorously test this unified model using three distinct retrieval strategies for ICL demonstrations, all of which yields a consistent linear distribution, as illustrated in Figure 2c. This consistent alignment, with a Spearman correlation coefficient of 0.8979 and a $p$-value approaching zero between calculated power in Eq. 3 and practical accuracy, provides strong empirical evidence for the robustness of our theoretical model. These results confirm that both ICL and CoT not only can be calculated independently but are inherently compatible within the same theoretical structure, offering a powerful new way to describe and predict the behavior of LLM systems.
3.4 Extension Experiments on Various LLMs and Tasks.
To demonstrate the generality of our theoretical ECM, we conduct extension experiments across LLMs configured with varying power supply voltages to represent different capacities. Figure 2d reveals a consistent linear correlation over 0.8348 between theoretical output power estimates and empirical accuracy, confirming the ECM’s robustness in predicting performance. Beyond traditional mathematical reasoning tasks, we test ECM on diverse real-world problems, including multi-hop knowledge reasoning [44], medical probing [7], and robotic planning [34]. As shown in Figure 2e, ECM achieves Spearman coefficients exceeding 0.86 and R2 exceeding 0.77 across all tasks, underscoring its versatility in both cognitive and practical applications.
4 Explanation based on ECM
In this section, we mainly pay attention to how to use ECM to explain the effectiveness of existing prompting methods.
4.1 Explanation for ICL Strategies
Enhanced embeddings facilitate accurate semantic magnetic field strength estimation and retrieval.
Recent advancements in embedding models have shown that improved text representations directly enhance LLM’s efficacy in the retrieval process [21]. This improvement is tied to the precise estimation of semantic magnetic field strength. As shown in Figure 3a, we employ three models—BERT [9], RoBERTa [22], and BGE [4]—to represent and retrieve demonstrations in ICL. Among these, BGE achieves the highest degree of alignment, with a Spearman correlation coefficient of 0.9106 between theoretical power and empirical accuracy, significantly surpassing the other models. This result highlights superior embedding ability attributed to the accurate semantic magnetic field estimation.
Figure 3: The Experimental Explanation for Existing ICL Strategies.
a,
The relationship between theoretical power and actual accuracy for various encoders is assessed. BGE achieves the highest Spearman correlation coefficient of 0.9106. b, A comparison of the Vanilla Encoder and the Meta-Recognition Representation reveals the latter’s ability to enhance the linear correlation between theoretical power and practical accuracy, from 0.9130 to 0.9337. c, The distribution of magnetic field strengths in generated demonstrations (with CDF representing cumulative distribution function and CCDF representing complementary cumulative distribution function). Further, we demonstrates the effectiveness of different Synthetic-CoT with degree of different degree of semantic field strength. d, The effectiveness of Synthetic-CoT is analyzed across three conditions: Positive Synthetic-CoT (demonstrations with positive extra voltage from ICL based on positive magnetic fields), Zero-shot CoT (no demonstrations), and Negative Synthetic-CoT (demonstrations with negative extra voltage from ICL based on negative magnetic fields). The left scatter plot displays the performance distribution of these conditions on the power-accuracy landscape. The right bar chart highlights factors contributing to Synthetic-CoT’s optimization of extra voltage provided by ICL.
Meta Recognition Enhance Magnetic Field Strength Estimation Modeling.
The study of Didolkar et al. [10] introduces Meta-Recognition, a method leveraging GPT-4 [1] to decompose complex problems into modular sub-capabilities for ICL representation, improving LLMs’ performance. We attribute its effectiveness to the enhanced sample representation modeling, which enhances ICL voltage and subsequently boosts output power and accuracy. As illustrated in Figure 3b, Meta-Recognition significantly increases the correlation between theoretical power and practical accuracy, improving from 0.9130 to 9337. It underscores that representation based on sub-capability decomposition effectively enhances the precision of magnetic field strength modeling.
Synthetic-CoT predominantly generates samples with positive magnetic fields, enhancing model performance.
Synthetic-CoT, a widely used approach in ICL, enables LLMs to autonomously generate ICL demonstrations, thereby improving performance [33]. As illustrated in Figure 3d, all ICL demonstrations produced by Synthetic-CoT exhibit positive semantic magnetic field contributions, which are strongly correlated with enhanced model accuracy. This effect is likely due to the positive voltage induced by the magnetic field, which optimizes circuit functionality. As shown in Figure 3d, the larger ICL magnetic field strength corresponds to the larger power and the higher accuracy. To further investigate the impact of magnetic field polarity on model performance, we analyzed demonstrations with negative magnetic field. Including such examples led to a 21.47% reduction in accuracy and a substantial decline in theoretical power output (Figure 3d). These findings underscore the critical role of positive magnetic field demonstrations in the superior performance achieved by Synthetic-CoT.
Figure 4: The Experimental Explanation for Existing CoT Strategies.
a,
The distributions of accuracy, power, and resistance for Zero-shot without CoT and Zero-shot with CoT are presented. The average resistance for Zero-shot without CoT is 10.93, significantly higher than the 1.50 observed for Zero-shot with CoT. b, Mechanistic analysis of Tool Usage (TU) and Program-of-Thought (PoT) within the ECM. Initially, TU, modeled on the original CoT, had non-zero computational resistance ($R_{C}$), limiting power-accuracy correlation. Assuming $R_{C}=0$ achieves a 10% improvement in the Spearman coefficient. PoT, representing TU with structured code, further enhanced performance by keeping $R_{C}=0$ and reducing planning resistance $R_{P}$, achieving another 10% improvement in power-accuracy correlation. c, A circuit-based rationale to explain how to utilize self-consistency and coverage to enhance performance. Self-consistency reduces total resistance by processing multiple reasoning resistors in parallel (sampling $n$ solutions and selecting the output with the highest consensus). In contrast, coverage minimizes resistance by increasing sample size with negligible resistance of voting resistor (sampling $n$ instances and passing the task if any succeeds).
4.2 Explanation for CoT Strategies
Direct Answering Introduces Greater Logical Resistor Compared to CoT
Recent studies suggest that generating direct answers without intermediary reasoning steps increases cognitive load on the model [13, 17]. We hypothesize that the absence of explicit reasoning pathways introduces significant resistance of “logical resistor,” leading to weaker performance compared to Zero-shot CoT. To test this, we analyzed performance and power correlations for direct answering and Zero-shot CoT on a power-accuracy graph, finding a strong linear correlation of 0.9131 ($p$-value $<0.01$). The average resistance for direct answering is calculated as 10.93, substantially higher than the 1.50 observed for Zero-shot CoT. This discrepancy corresponds to a performance reduction of over 30% for direct answering. These findings indicate that the lack of a CoT framework imposes additional greater logical resistor, hindering the model’s reasoning capabilities and resulting in diminished performance.
Tool Usage optimizes computational resistor, and PoT further reduces planning resistor, improving model performance.
Chen et al. [5] propose that Program of Thought (PoT) and Tool Usage (TU) enhance computational reasoning capabilities to a theoretical maximum, effectively minimizing reasoning complexity to zero, akin to eliminating the resistance to zero. Inspired by this, we investigate the correlation between CoT and TU models under this "zero-resistance" assumption. Our analysis revealed that modeling TU as a circuit with zero resistance of computational resistor increased the Power-accuracy Spearman correlation by over 10% compared to the original model, as illustrated in Figure 4b, which validates our hypothesis.
Furthermore, we hypothesize that PoT, leveraging programming languages, reduces planning resistance. As shown in Figure 4b, modeling PoT as a circuit with lower planning resistance achieved a correlation score exceeding 0.9423 between theoretical power and accuracy—significantly higher than TU. This finding highlights the superior effectiveness of PoT in improving model performance.
Self-consistency and Inference Scaling Law via Parallel Resistor in Circuits
Self-consistency, as demonstrated by Wang et al. [40], utilizes multiple reasoning paths simultaneously, and determines final decision through majority voting, enhancing performance robustness and accuracy. Here, we interpret its effectiveness using a circuit model, which introduce parallelization of original resistors ($R_{\text{CoT}}$), significantly reducing the total system resistance. As shown in Figure 4c, the practical accuracy of the self-consistency strategy aligns closely with the theoretical output power curve, validating the correctness of our proposed model. Recent findings on the Inference Scaling Law [43, 28] further demonstrate that increasing the number of parallel samples, $n$, substantially improves accuracy. As $n$ grows, the resistance value of CoT resistor approaches zero, optimizing the reasoning process. Further, when coverage introduces an absolutely correct answer verification process with zero resistance of verification resistor, exceeding the gains achieved by self-consistency alone. The alignment between theoretical predictions and empirical results, also evident in Figure 4c, reinforces the validity of this model. These results highlight the central role of self-consistency and inference scaling in enhancing the performance of circuits.
Figure 5: The Experimental Exploration for ICL Strategies.
a,
The relationship between the introduction of demonstrations and model performance. Using Reverse ICL + CoT, which introduces demonstrations with negative magnetic fields, impairs the model’s reasoning ability and results in performance falling below that of Zero-shot CoT. In contrast, Retrieval-ICL + CoT, incorporating demonstrations with positive magnetic fields, significantly improves performance. b, The upper part of the figure contrasts Diversity-oriented Static ICL (DS-ICL), where demonstrations are retrieved for the whole task, and Similarity-oriented Dynamic ICL (SD-ICL), where demonstrations are retrieved for each request. In SD-ICL, demonstrations are more aligned with test requests, creating a stronger semantic magnetic field. The lower part highlights Diversity-oriented Dynamic ICL (DD-ICL), where increasing sampling range enhances the diversity but decreases the magnetic field strength of retrieved demonstrations, resulting in a more diverse but weaker strength, leading to less accuracy. c, The left figure illustrates the demo-optimization process, a methodology designed to refine the demonstrations used in ICL. The right figure shows that demo-optimization outperforms traditional prompt optimization by improving theoretical power.
5 Exploration based on ECM
In this section, we primarily focus on how to use ECM to explore phenomena that have never been well explained and more novel prompting methods.
5.1 ICL Exploration
When Few-shot CoT Performs Worse than Zero-shot CoT?
Despite the growing importance of ICL and CoT, a critical question remains: under what conditions is ICL essential, and when does zero-shot CoT suffice? To investigate this, we systematically examine the interaction between these two strategies. As shown in Figure 5a, our findings reveal a phenomenon we term “reverse ICL," where demonstrations associated with a negative semantic magnetic field significantly impair the output power of circuits, thereby decreasing the model’s reasoning ability and resulting in worse performance compared to zero-shot CoT. This suggests that not all demonstrations contribute positively to the learning process. In contrast, demonstrations with a positive semantic magnetic field lead to an enhancement in performance. These results highlight the critical importance of selecting demonstrations based on their associated semantic magnetic field, which plays a crucial role in determining the success of ICL.
The influence of diversity in ICL retrieval: a static vs. dynamic perspective
In ICL retrieval, two paradigms are commonly studied: static ICL, where the set of demonstrations remains fixed during testing, and dynamic ICL, where demonstrations are selected in real-time based on the specific test sample. Previous studies [19] reveal that semantic diversity in static ICL enhances performance by exposing the model to a broader range of contextual cues. Conversely, in dynamic ICL, diversity does not provide the same benefits and may even hinder performance [31]. This raises the question: Why does diversity yield such contrasting outcomes in these retrieval modes? To this end, we hypothesize that the benefits of diversity in static ICL can be likened to a magnetic field. Diverse demonstrations create a robust and expansive field, offering the model access to varied directional signals as test samples change. This broad-spectrum magnetic field stabilizes the model’s decision-making process, improving its generalizability and overall performance. As illustrated in Figure 5b, static ICL benefits from this diversity, which amplifies the model’s predictive power compared to random selection strategies. In dynamic ICL retrieval, however, the focus shifts toward selecting demonstrations closely aligned with the test sample. This approach optimizes the magnetic field for stronger, more targeted signals in the relevant direction. In this context, diversity becomes minimally critical; the alignment and intensity of retrieved demonstrations take precedence. As shown lower panel in Figure 5b, diverse but less aligned demonstrations enable less accuracy.
Demo-Optimization: a novel approach to enhance demonstration efficacy in ICL
Recent advancements in prompt-based learning leverage feedback from outputs as model gradients, enabling large models to autonomously refine their prompts for more efficient learning [45]. Expanding on this foundation, we propose demo-optimization, a methodology tailored to optimize the demonstrations used in ICL. Demo-optimization focuses on balancing the relative effectiveness between user requests and demonstrations (See Method for implementation details). As illustrated in Figure 5c, demo-optimization significantly outperforms traditional prompt optimization approaches, which primarily refine instructions. This method not only improves accuracy but also theoretically strengthens reasoning capabilities. Moreover, it addresses the persistent underfitting problem in prior ICL methods, offering a more precise and scalable strategy for enhancing demonstration-driven learning.
Figure 6: Experimental Evaluation of Chain-of-Thought Strategies and Power-Optimized Performance on Complex Tasks
a,
Overview of the fine-grained self-consistency strategy in circuits. By decomposing parallel CoT resistors into smaller units, this approach effectively reduces overall circuit resistance value. b, Overview of the Chain-of-Verfication (CoV) strategy in circuits. In addition to optimizing parallel CoT resistances, this method incorporates the concept of voting resistors in self-consistency to further reduce overall circuit resistance. The CoV strategy achieves this by sampling multiple voting results and validating them to finalize the decision. c, Leveraging stronger magnetic fields and reducing logical resistance significantly enhances circuit performance. The power-enhanced model achieved notable results in the IOI (left figure) and the IMO (right figure). d, The methodology for incorporating sufficient theoretical power into manual ratings during idea generation scenarios. This approach demonstrates a significant improvement over the original solution by enhancing the accuracy and reliability of the evaluation process.
5.2 CoT Exploration
Optimize the same proportion of capabilities. The greater the original difficulty of the task corresponding to the capability, the greater the impact on the final performance.
We can prove that optimizing a resistor with higher resistance to its original value $\frac{1}{k}$ results in a greater reduction in overall resistance of reasoning resistor $R_{\text{CoT}}$, yielding a more significant performance improvement compared to optimizing a lower resistance (See in Theoretical Proof). As shown in the meta-analysis in Appendix A, when solving questions that demand advanced planning, o1 [28]’s reasoning method, which emphasizes extensive exploration, significantly enhances the most challenging reasoning components. This improvement ultimately boosts overall performance on complex reasoning tasks. However, for tasks where the hardest component requires extensive domain-specific knowledge, insufficient optimization of such knowledge limits performance gains in those domains.
Fine-grained Self-consistency achieves better performance
Easily, we can prove that, with a stable supply voltage, we can split a self-consistency ensemble into more fine-grained step-by-step self-consistency by conducting majority voting verification for each step, so that each resistance value of verification resistor can be smaller, and we can obtain larger output power (see Theoretical Proof). Indeed, as the number of self-consistency steps increases, the theoretical power value gradually increases. As shown in Figure 6a, the corresponding performance also gradually increases. At the same time, the corresponding upper bound and lower bound also satisfy the original self-consistency and coverage. This not only proves the effectiveness of our proposed fine-grained self-consistency method but also proves the practical usability of the theoretical model.
Chain-of-Verification enhances model performance with parallelized verification resistors
In conventional self-consistency frameworks, verification is typically conducted through series resistors. We propose an innovative approach, referred to as the “chain-of-verification,” which involves two key looped stages: first, verifying the correctness of individual outputs in parallel; and second, merging these verified results and subsequently meta-validating the correct to confirm the overall result of the merged outputs. It can enhance model performance by reducing the overall verification resistance. As depicted in Figure 6b, parallelization initially boosts performance by distributing verification tasks and lowering verification resistance. However, this improvement has limits. As parallelization increases, the overall resistors for parallel verifiers begins to rise. Beyond a critical threshold, this combined resistor surpasses the original self-consistency resistor, reversing the performance gains. As depicted in Figure 6b, this also demonstrates excessive chain-of-verification introduces diminishing returns. Thus, the chain-of-verification framework underscores the importance of balancing parallelization and resistor to optimize model verification efficiency.
6 Best Practice
To evaluate the impact of sufficient power on LLM performance, we analyzed its effects across three domains: the International Olympiad in Informatics (IOI), the International Mathematical Olympiad (IMO), and Natural Language Processing (NLP) idea generation using the AI-Scientist framework [23], by ensembling strategies outlined in Exploration Section. Our findings, shown in Figure 6c, demonstrate that augmenting the o1-preview model with enhanced power results in significant improvements. The power-enhanced o1-preview outperformed AlphaCode [20] and other state-of-the-art models on the IOI-level test set (code_contests [20]). Crucially, this boost consistently surpassed configurations lacking sufficient power, underscoring its pivotal role in model efficacy. In competitive contexts, such as IOI 2024 and IMO 2024, the powered o1-preview outperformed nearly 80% of global participants, earning silver and bronze medals. This highlights power as a decisive factor in high-stakes scenarios where traditional optimizations often fall short. Beyond correctness-focused tasks, as shown in Figure 6d, sufficient power also enhanced performance in creative challenges. For instance, in NLP-related idea-generation, it facilitates the better generation of novel and impactful research ideas (over 1 point of overall score improvement on average), expanding the potential of LLMs in both objective and subjective tasks.
7 Conclusion & Discussion
The Electronic Circuit Model (ECM) offers a unified framework for understanding and optimizing the ICL and CoT mechanisms, marking a significant advancement in LLM research. This framework elucidates model behavior while providing a quantifiable approach to enhance LLM performance. Our experiments demonstrate that ECM effectively captures model intricacies, serving as a reliable tool for analyzing and improving AI capabilities. Moreover, applying power-based optimization derived from ECM to advanced reasoning challenges, such as the International Olympiad in Informatics and the International Mathematical Olympiad, yielded superior results compared to leading competitors. These outcomes validate the theoretical foundation of ECM and highlight its transformative potential in improving AI reasoning and learning efficiency, with broad implications for next-generation intelligent systems.
8 Code Availability
The code and relevant data are available at https://github.com/LightChen233/ECM.
A Method Implementation Details
A.1 Theoretical Model
Faraday’s Law of In-Context Learning
In-context learning (ICL) can be likened to the behavior of a semantic magnetic field generated by a sub-power source, as described by Faraday’s Law. The positive direction of the semantic magnetic field is defined by the query semantic vector $\overrightarrow{S}_{q}$. Given $N$ context samples represented by semantic vectors $\overrightarrow{S}_{i}$, the total semantic magnetic field strength $\Phi^{B}_{0}$ experienced by the model is the projection of these vectors onto $\overrightarrow{S}_{q}$. Mathematically:
$\Phi^{B}_{0}=\sum^{N}_{i=1}\cos\theta_{qi}\cdot\left\lvert\overrightarrow{S}_{i}\right\rvert=\sum^{N}_{i=1}\frac{\overrightarrow{S}_{q}\cdot\overrightarrow{S}_{i}}{\left\lvert\overrightarrow{S}_{q}\right\rvert}$ (4)
where $\cos\theta_{qi}$ represents the angle between $\overrightarrow{S}_{q}$ and $\overrightarrow{S}_{i}$.
As new context samples are processed, the semantic field strength $\Phi^{B}_{0}$ is hypothesized to decay linearly over time:
$\Phi^{B}(t)=-\lambda\Phi^{B}_{0}t,$ (5)
where $\lambda$ is the decay constant representing the rate at which the model forgets or absorbs information. According to Faraday’s Law of electromagnetic induction, the rate of change in the semantic magnetic field induces an electromotive force, which drives effective ICL. With a time step $\Delta t$ for each reasoning step, the induced EMF, $\mathcal{E}_{\text{ICL}}$, is:
$\mathcal{E}_{\text{ICL}}=-\frac{\Delta\Phi^{B}(t)}{\Delta t}=\lambda\Phi^{B}_{0}.$ (6)
Substituting $\Phi^{B}_{0}$ into this equation yields:
$\mathcal{E}_{\text{ICL}}=\lambda\sum^{N}_{i=1}\frac{\overrightarrow{S}_{q}\cdot\overrightarrow{S}_{i}}{\left\lvert\overrightarrow{S}_{q}\right\rvert}.$ (7)
In conclusion, this analogy offers a structured framework to analyze how information flows during ICL. By extending the dynamics of ICL through the principles of electromagnetic induction, we provide a mathematical model to explain how models incorporate and gradually forget information as they process contextual examples.
Ohm’s Law of Chain-of-Thought
In this framework, each sub-difficulty of a model introduces a distinct level of difficulty, analogous to resistors in series. Formally, following Chen et al. [5], the ease with which a model can solve a particular task is represented by its reasoning boundary. Let $\mathcal{B}_{\text{CoT}}$ denote the overall reasoning boundary of a Chain-of-Thought process, and $\mathcal{B}_{i}$ denote the reasoning boundary of the $i$-th sub-capability. The overall reasoning boundary can be expressed by the harmonic sum:
$\mathcal{B}_{\text{CoT}}=\frac{1}{\sum^{K}_{i=1}\frac{1}{\mathcal{B}_{i}}},$ (8)
where $K$ denotes the number of sub-capabilities. Reasoning resistor, $R$, quantifies the difficulty of a reasoning process and is mathematically defined as the reciprocal of the reasoning boundary. A higher boundary corresponds to an easier problem and lower resistor. Specifically, we define the sub-difficulty of the $i$-th sub-task as:
$R_{i}=\frac{1}{\mathcal{B}_{i}}.$ (9)
Substituting this into Equation 8, the total reasoning resistor for the CoT process is expressed as:
$R_{\text{CoT}}=\sum^{K}_{i=1}R_{i},$ (10)
where $R_{\text{CoT}}$ represents the combined difficulty, analogous to total resistor in a series circuit. Overall, the CoT reasoning framework is further formalized by the Ohm’s law analogy:
$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}}{R_{\text{CoT}}+R_{0}},$ (11)
where $R_{0}$ denotes a output resistor with static resistance value. This modular formulation effectively quantifies cognitive task complexity through combined reasoning resistor.
A.2 Genral Experimental Settings
All experiments, unless stated otherwise, are conducted on GPT-3.5-Turbo. Following Wei et al. [41], CoT experiments select 5 manually constructed demonstrations for all multi-step reasoning tasks. Top-p is selected from $\{0.95,1\}$, and Temperature is selected in range $[0,1]$ for robustness evaluation. Experiments are primarily conducted on the BigGSM dataset [5], a benchmark for mathematical reasoning. In ICL, demonstrations are dynamically retrieved using a 5-shot approach from the GSM8K development dataset [8] by default. To formalize the demonstration selection process, we retrieve the semantic magnetic field strength $\Phi^{B}_{0}$ by default, as defined in Eq. 4. All demonstration representation methods employ the meta-recognition [10] framework by default: GPT-4o [1] extracts sub-capability tags, and Roberta [22] concatenates and represents these tags in the hidden space. In addition, only the magnetic field strength $\Phi^{B}_{0}$ is calculated and the resistor value $R_{\text{CoT}}$ directly adopts the fitting values from Chen et al. [5]. In addition, other parameters are fitted values using the verification set. Different models fit different voltages $\mathcal{E}_{model}$. The whole experiment shares a $R_{0}$ value. Different demonstration representation methods have different $\lambda$. All experiments involving variance were sampled 3 times at different temperatures.
A.3 Model Verification Experiment Details
CoT Ohm’s Law Verification
To validate the Ohm’s Law assumption in the CoT approach, we conduct the following experiment. To eliminate potential interference from ICL and ensure effective stimulation of CoT capabilities, we append the phrase “Let’s think step-by-step!" to the user query. This adjustment aims to stimulate and isolate the CoT mechanism and prevent ICL-related influences. We then perform sampling based on power levels from the BigGSM dataset, using 30-unit intervals to map calculated power to accuracy [^1] The absolute power values are meaningless due to inconsistencies in semantic space scaling across encoding methods. However, their relative values within a method correlate positively with real accuracy.. To reduce random errors and ensure sufficient sampling points, we retain data points with at least 10 samples. Accuracy is calculated for each sampling interval to analyze the relationship between power levels and model performance, enabling an effective evaluation of the correlation between sampling power and accuracy.
ICL Faraday’s Law Verification
To validate the semantic magnetic field hypothesis in ICL, we observe that without CoT reasoning, performance on complex mathematical tasks often declines to near-zero levels, thereby cannot easily observe their relevance. Thus, this experiment incorporates both CoT reasoning and ICL. While the CoT reasoning format remains constant, we focus on varying the similarity measures used for sample retrieval through ICL. The central adjustment involves applying different similarity measures for sample retrieval, directly affecting the model’s performance on complex tasks. These measures also help calculate the additional “voltage" introduced by semantic magnetic fields, quantifying their influence on performance. All demonstration representation methods employ BGE [4] to represent the request in the hidden space.
Unified circuit mechanism verification
To further validate the effectiveness of our unified framework for modeling ICL and CoT, we conduct experiments integrating both paradigms into different retrieval settings. These experiments test the robustness and rationality of the proposed theoretical framework. Specifically, we examine three retrieval methods: random retrieval, forward magnetic field retrieval, and reverse magnetic field retrieval. In the random retrieval setup, samples are selected randomly from the development set as ICL demonstrations for each test sample. The Retrieval-ICL selects samples with maximized magnetic field intensity in the corresponding test sample’s direction. Conversely, the Reverse-ICL focuses on samples with minimal (even negative) magnetic field intensity, enabling exploration of contrasting ICL examples. This comprehensive validation highlights the robustness of our framework across diverse retrieval strategies and empirically demonstrates its generalizability in both ICL and CoT scenarios.
Extension Experiment Details on Various Models and Tasks
To further enhance the applicability of ECM, we conducted experiments across a series of models, where each model was designed to fit specific baseline voltage values, $\Phi^{B}_{0}$, and compute the corresponding power output. Subsequently, all models were consolidated onto a single graph, enabling us to analyze the correlation between output power and relative accuracy. As illustrated in Figure x, the results reveal clear and detailed correlations across the models. However, when accuracy exceeds 90%, the scarcity of samples requiring sufficient power in this range, likely due to limitations in the dataset, leads to poor fitting in this region.
Furthermore, our experiments on HotpotQA and Medical Probing closely adhered to the experimental settings established by Chen et al. [5]. However, since Medical Probing includes rationale annotations limited to background knowledge rather than chain reasoning, we supplemented portions of the reasoning process to facilitate a more precise correlation analysis. Additionally, the data for robotic planning was directly extracted from the running logs provided by Valmeekam et al. [34].
A.4 Model Explanation Experiments
A.4.1 ICL Explanation
Embedder Comparison
To evaluate how different embedders influence model performance, we replace the sample encoder with BERT [9], RoBERTa [22], and BGE [4]. These embeddings are used for sample retrieval in ICL to minimize potential biases that may arise if a single embedder disproportionately emphasizes specific data features. For instance, an embedder might favor samples aligned with its encoding patterns, leading to distortions in the relationship between retrieved representations and model performance. By employing diverse embedders, we aim to mitigate such biases and promote more accurate theoretical modeling for ICL.
Synthetic-CoT Implementation
This section examines the implementation of Synthetic-CoT within the ECM framework to assess its impact on reasoning performance. Synthetic-CoT autonomously generates in-context demonstrations with chain-of-thought reasoning paths. We hypothesize that these demonstrations align the system’s magnetic field, reducing logical resistor and enhancing overall circuit power output. To validate this, we generate a series of demonstrations with two categories by Synthetic-CoT, including:
  • Positive Magnetic Field Demonstrations: In-context demonstrations generated by Synthetic-CoT are positively aligned with the magnetic polarity of user queries direction.
  • Negative Magnetic Field Demonstrations: In-context demonstrations generated by Synthetic-CoT are with opposing magnetic polarity of user queries direction.
We analyze the magnetic field contributions of these categories, focusing on their differing effects. Subsequently, we integrate the computed power and measured accuracy to further investigate their operational mechanisms.
A.4.2 CoT Explanation
Direct answer output
We analyze the effect of direct answer output on reasoning performance through the ECM framework. This approach bypasses intermediary reasoning steps by appending the directive, “Let’s directly output the result without additional reasoning content,” which generates significant logical resistor within the system. To quantify this, we gradually improve the resistance value of reasoning resistor $R_{\text{CoT}}$, which meets largest Spearman correlations in the validation set, and employ the $R_{\text{CoT}}$ to the test set.
Tool-Usage and PoT
We further investigate how Tool-Usage and Program-of-Thought (PoT) mitigate specific resistors in the reasoning process, focusing on their impact under the ECM framework on computational and planning resistors:
  • Tool-Usage: Tool-Usage aims to reduce computational resistor by offloading complex calculations to external tools. Modeled as a zero-resistor computational circuit in our experiments, Tool-Usage enhances the Power-Accuracy Spearman correlation.
  • Program-of-Thought (PoT): PoT targets planning resistor by employing structured programming logic to enable clearer and more efficient reasoning. This advantage stems from PoT’s superior ability to reduce planning resistor, facilitating finer-grained logical alignment. [^2] All implementation prompts come from Chen et al. [5]
These findings confirm that while Tool-Usage primarily addresses single-step computational resistor, PoT excels in optimizing overall planning resistor. Both methods significantly enhance performance within the ECM framework, highlighting their complementary contributions to reducing resistor and improving reasoning capabilities.
Self-consistency and Coverage Design
This experiment evaluated the self-consistency strategy proposed by Wang et al. [40], which employs multiple parallel reasoning paths and a majority-vote mechanism for the final decision. This strategy is hypothesized to enhance performance by reducing the system’s resistance to diverse inputs, analogous to parallel resistors in a circuit, where total resistor decreases compared to series connections. By applying this analogy, we demonstrate how parallel reasoning can lower cognitive load, improving accuracy and efficiency in decision-making.
To examine the influence of inference scaling on model performance, we investigated the relationship between the number of parallel samples, $k$, and the metric like Accuracy or Pass@k, as described in the Inference Scaling Law. Specifically, Pass@k measures the accuracy of Coverage method which obtains a correct result when at least one sampled prediction is accurate. We modeled the reasoning process as a resistive circuit, where each sample represents a parallel resistor. As $k$ increases, the resistance of reasoning resistor approaches zero, reflecting a refined reasoning process and enhanced accuracy. Further, Coverage methods introduce zero verification resistance introduce much less overall resistance values, larger output power, and higher related performance. The experiment systematically varied $k$ from 1 to 100 in increments of 5, recording the accuracy and Pass@k metric under each condition [^3] We remove those generated samples are totally the same to keep the independent sampling.. We track individual reasoning path resistance, comparing these values with observed accuracy to validate the circuit analogy. The effectiveness of Self-consistency and Coverage methods can be mathematically proved based on the ECM detailed in Theoretical Proof.
A.5 Model Exploration Experiments
A.5.1 ICL Exploration
Diversity
This study examines the differing impacts of semantic diversity on static and dynamic ICL retrieval. We map all samples into the hidden space using RoBERTa. In static ICL, we employ a fixed set of demonstrations with the highest semantic diversity across the task, hypothesizing that this diversity enhances model performance by providing a wider range of contextual cues. In contrast, dynamic ICL selects demonstrations in real-time based on the test sample’s characteristics, prioritizing semantically similar examples that align closely with the input. To test these, we evaluate both retrieval strategies. For static ICL, we compare models exposed to high-diversity and random demonstration sets, measuring performance using accuracy. For dynamic ICL, we contrast random selection with retrieval strategies emphasizing semantic alignment. We hypothesize that semantic diversity improves static ICL by stabilizing decision-making, while in dynamic ICL, retrieving highly relevant demonstrations yields better accuracy. In addition, in order to prove that diversity is ineffective in the setting of dynamic ICL, we sort the semantic magnetic field strength and select as diverse samples as possible among the top-$k$ samples as ICL demonstrations.
Demo-Optimization
This experiment introduces demo-optimization, a novel method for optimizing demonstrations in ICL. The approach dynamically refines demonstrations based on feedback from the LLM’s outputs, which serve as gradients for improvement. Specifically, initial ICL demonstrations in prompting are selected, with both development dataset for optimization and optimization prompt configurations prepared. A feedback loop guides adjustments to the demonstrations, improving their clarity, relevance, and alignment with prompt requirements. Demonstrations are iteratively refined, and evaluated for contextual alignment with user requests in the development dataset. The method is compared against static and whole-prompt-optimized baselines, with performance assessed using task accuracy, output consistency, and learning efficiency.
A.5.2 CoT Exploration
CoT Optimization Proportion
Figure 7: Meta-analysis for the model improvements on different tasks. The data comes from Wu et al. [42], Valmeekam et al. [35] and OpenAI [28]. We collect evaluation data related to plan and knowledge based on OpenAI [28]. In addition, we remove data points with final results exceeding 90%, because these data have exceeded human levels, resulting in potential labeling bias and test set labeling errors.
To evaluate the improvement in reasoning resistance $R_{i}$ to the target level of $\frac{1}{k}$, which depends heavily on the task’s reasoning demands, we conducted a meta-analysis of experimental results from diverse datasets. These datasets are divided into two main categories: (1) tasks requiring advanced planning and logical reasoning, such as mathematical derivations and multi-hop question answering, and (2) tasks reliant on extensive domain-specific knowledge, such as chemistry and physical question answering. As shown in Figure 7, our meta-analysis demonstrates that tasks requiring advanced planning gain the most from enhanced exploratory reasoning, significantly improving challenging components (e.g., a 2.8-fold optimization in planning capability) and overall task performance (e.g., a 3.6-fold optimization on AIME, a benchmark emphasizing long-term planning). In contrast, tasks relying heavily on domain-specific knowledge exhibit limited improvement without targeted optimization in those areas. This underscores the importance of domain-aware reasoning strategies. For example, tasks involving complex reasoning and knowledge integration, such as GPQA (chemistry, physics) and HotpotQA, show performance improvements directly proportional to advancements in the related knowledge domains (approximately 1:1 ratio). The effectiveness and the underlying reasons for such phenomena can be mathematically analyzed and proved based on the ECM detailed in Theoretical Proof.
Fine-grained Self-consistency
In this experiment, we examine how decomposing a self-consistency ensemble into smaller, incremental majority-voting steps reduces reasoning resistor in circuits, compared to the conventional single-step approach. The circuit operates at a stable supply voltage of 5 units, and reasoning resistance value and output power are evaluated as the number of steps increases. Each majority-voting step simplifies logical complexity and reduces overall resistance by incorporating less verification resistance in the decision-making process. Our findings demonstrate a consistent decline in reasoning resistance alongside a rise in output power, confirming that finer-grained majority-voting processes improve both efficiency and power output in circuits. The effectiveness of this method can be mathematically proved based on the ECM detailed in Theoretical Proof.
Chain-of-Verification
Traditional self-consistency frameworks rely on a sequential verification model, where each step depends on the completion of the previous one. This structure often incurs high computational costs, as subsequent tasks cannot begin until earlier ones are resolved. To overcome this limitation, we introduce the “chain-of-verification (CoV)," a novel strategy that parallelizes verification tasks into independent components. Each component operates concurrently, validating distinct parts of the system without interdependence. This parallel approach reduces overall computational resistance by distributing the workload across multiple verifiers, thereby enabling more effective validation. As illustrated in Figure 6b, parallelization not only accelerates the verification process but also enhances model performance by mitigating bottlenecks. This improvement is particularly pronounced in large-scale systems, where sequential verification would otherwise impede efficiency. The effectiveness of CoV can also be mathematically proved based on the ECM detailed in Theoretical Proof.
A.6 Best Practice
In practice, we integrate all proposed strategies into the O1-preview model to address IOI, IMO problem-solving, and idea-generation tasks for much less resistance value of planning resistor.
IOI Task
For the IOI task, we employ the following methods: First, using O1-preview, we maximize the optimization of the resistance value of its “planning resistor." Additionally, five high-intensity semantic field problems from the code_contest training dataset are retrieved as demonstrations, boosting the voltage supplied by additional ICL fields. Second, we apply the Coverage technique, sampling 100 instances and leveraging both provided and model-generated test cases in parallel, effectively reducing resistance in chain-of-thought reasoning. This process, executed in parallel, effectively reduces the resistance in the chain-of-thought reasoning. Furthermore, we incorporate the CoV concept to iteratively validate the final outputs, thereby optimizing the final performance. Additionally, the CoV framework is employed to iteratively validate outputs and optimize overall performance. To prevent data leakage from competition datasets in C++, competition code is generated in Python. Manual adjustments ensure compatibility between Python and C++ input/output handling, preserving the core logic for fair comparative testing.
IMO Task
A similar approach is adopted for the IMO task. O1-preview is optimized for planning resistor, and five high-intensity semantic field problems from past IMO datasets are retrieved as demonstrations to enhance ICL field voltage. The Coverage technique is again employed, sampling 100 instances in parallel to reduce reasoning resistance. We introduce the fine-grained self-consistency approach within the CoV framework, enabling iterative validation of outputs and achieving notable performance gains. Additionally, we only evaluate answers and do not consider process scores.
Idea-Generation Task
We utilize the AI Scientist framework [23] to generate over 10 ideas per domain and draft academic papers, limited to the abstract and introduction sections to address ethical concerns by avoiding experimental claims. A 16-member review panel, including 2 Superior reviewers (6+ years of experience), 8 Senior reviewers (4+ years), and 6 Junior reviewers (2+ years), ensures reliability. Each paper is reviewed by at least four reviewers, including two Senior or Superior reviewers, following NeurIPS guidelines. High-quality examples with strong semantic fields, sourced from past ICLR, NeurIPS, and CoLM openreview data, are used to guide generation. The Coverage technique, sampling of 100 instances, and iterative validation through the CoV framework further refine the process.
Figure 8: Neuron activation heatmaps for the most activated attention layer under three reasoning strategies: (a) Zero-shot CoT, (b) ICL with Semantic Magnetic Field Retrieval, and (c) Program of Thought (PoT).
A.7 Internal Mechanism underlying the ECM
To investigate the internal mechanisms underlying ECM, we conduct experiments in the LLaMA3-8B [12] focusing on three strategies: Zero-shot CoT, ICL with Semantic Magnetic Field Retrieval, and PoT. By analyzing neuron activation values and frequencies—particularly in the most activated attention layer—we aim to elucidate the relationship between reasoning strategies and neural dynamics. These findings support the theoretical framework of ECM. The experiments use the BigGSM dataset, a standard benchmark for mathematical reasoning tasks. For ICL, demonstrations are dynamically retrieved using the BGE model in a 3-shot configuration. Figure 8 shows neuron activation heatmaps for the most activated attention layer under each reasoning strategy. For Zero-shot CoT, the heatmap reveals sparse activations, with limited regions of intense activity, highlighting the difficulty of reasoning without contextual support. Conversely, ICL and PoT display more distributed and elevated activation patterns. These suggest that the underlying reasoning mechanisms engage a broader spectrum of neurons, with comparable activation dynamics reflecting similar reasoning power across these strategies.
B Theoretical Proof
B.1 Analysis of Effectiveness of Self-consistency and Coverage Design
B.1.1 Analysis of Effectiveness of Self-consistency
First, according to the series resistor assumption of CoT, the total resistance of series resistors can be expressed as:
$R_{\text{all}}=R_{0}+R_{\text{CoT}}.$ (12)
Building upon the illustration in Figure 4c, we hypothesize that self-consistency, achieved through multiple parallel reasoning processes and subsequent result aggregation by majority voting, can be represented as an aggregation resistor integrated alongside several parallel CoT resistors. In this framework, the effective resistance of the reasoning process transitions from $R_{\text{CoT}}$ to $\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}$, accompanied with a majority-voting resistor $R_{S}$ to determine the final outcome. Consequently, the total resistance associated with self-consistency, denoted as $R^{{}^{\prime}}_{\text{all}}$, is expressed as:
$R^{{}^{\prime}}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}},$ (13)
where $n$ is the number of samples, $R_{\text{CoT}}^{i}$ denotes the resistance corresponding to the $i-$th sample and $R_{S}$ accounts for the resistor due to self-consistency integration. This framework allows for rapid computation of the output power:
$P^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}})^{2}}.$ (14)
As the number of samples increases, the curves in Figure 4c reveal that the real performance curve is almost exactly consistent with the theoretical range. Furthermore, assume that $\{R_{\text{CoT}}^{i}\}_{i=1}^{\infty}$ is bounded thus we can easily prove that the overall resistance has a lower bound:
$\lim_{n\rightarrow+\infty}R^{\prime}_{\text{all}}=R_{0}+R_{S}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}=R_{0}+R_{S}.$ (15)
The corresponding performance bound is:
$\hat{P}^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{0}+R_{S})^{2}}$ (16)
As shown in Figure 4c, model performance aligns with this theoretical upper bound as sample size increases.
B.1.2 Analysis of Effectiveness of Coverage Design
Similarly, we hypothesize that coverage can be considered correct if at least one of multiple parallel reasoning is accurate. Under this assumption, the integration difficulty in self-consistency verification approaches zero. Analogously, this can be modeled as a zero-value combined resistance in parallel with multiple CoT resistors, where the coverage resistance of self-consistency $\hat{R}_{\text{all}}$ is given by:
$\hat{R}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}=R_{0}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}$ (17)
where $n$ denotes the number of samples. Using this, the theoretical output power can be similarly calculated as:
$P^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(\frac{1}{\sum_{n}\frac{1}{R_{\text{CoT}}}}+R_{0})^{2}},$ (18)
which are much more smaller than $P^{\text{self}}_{\text{out}}$ in the same $n$ sample size with $R_{S}\gg 0$. As the number of samples increases, the curve in Figure 4c demonstrates that the real performance aligns closely with the theoretical range. Building upon the same bounded assumption of $\{R_{\text{CoT}}^{i}\}_{i=1}^{\infty}$, this setup also defines a resistance lower bound:
$\lim_{n\rightarrow+\infty}\hat{R}_{\text{all}}=R_{0}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}=R_{0}$ (19)
and the corresponding lower-bound performance:
$\hat{P}^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}}{R_{0}}.$ (20)
As shown in Figure 4c, the model’s performance gradually converges toward this upper bound and much larger than the performance of self-consistency, demonstrating consistency with the theoretical predictions.
NOTE
According to the theoretical formulation in Eq. 18, in the reasoning phase where the resistance is significantly greater than $R_{0}$, the overall trend shows that Coverage is approximately proportional to the number of samples. Meanwhile, certain Inference Scaling Laws indicate that Coverage and the logarithm of the number of samples exhibit an approximately proportional relationship. These two observations are not contradictory, as our model assumes relatively independent sampling. However, we have observed that when the number of samples exceeds 100, an increasing proportion of the samples become duplicates rather than independent samples. Consequently, the probability of generating new samples follows the trend where $n$ samples are expected to yield approximately $\log n$ distinct reasoning samples.
B.2 Analysis of Effectiveness of Fine-grained Self-consistency
According to Eq. 13, the original circuit can be considered as a large parallel reasoning and verification process and the total resistor is as follows:
$R^{\prime}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}.$ (21)
As shown in Figure 6a, this process is decomposed into multiple sequential, smaller-scale parallel reasoning and verifications. This refinement transforms a single large parallel reasoning resistance with significant verification resistance into a sequence of smaller parallel reasoning resistance and smaller verification resistance. Formally, this is expressed as:
$R^{\text{fine}}_{\text{all}}=R_{0}+\sum_{j}R_{j}^{S}+\sum_{j}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{j}^{i}}},$ (22)
where $n$ denotes the sampling size, $R_{j}^{S}$ represents the small and fine-grained verification resistance for $j$-th step, and $R_{j}^{i}$ corresponds to smaller reasoning resistance for $j$-th step for $i$-th sampling. The total resistance can be broken down into resistance of sub-step resistor as $R^{i}_{\text{CoT}}=\sum_{j}R_{j}^{i}$.
We aim to demonstrate that our approach indeed optimizes the total resistance, i.e., $R^{\text{fine}}_{\text{all}}<R^{\prime}_{\text{all}}$. To prove this, we need to show that $R^{\text{fine}}_{\text{all}}-R^{\prime}_{\text{all}}<0$. The proof is as follows:
$R^{\text{fine}}_{\text{all}}-R^{\prime}_{\text{all}}$ $=\sum_{j}R_{j}^{S}-R_{S}+\frac{R_{\text{CoT}}}{n}-\sum_{j}\frac{R_{j}^{i}}{n}$ (23)
$=\sum_{j}R_{j}^{S}-R_{S}.$ (24)
It is evident that verifying correctness at the individual step level is simpler than verifying the correctness of the entire chain, so $\sum_{i}R_{i}^{S}<R_{S}$. Therefore, $R^{\text{fine}}_{\text{all}}-R^{\prime}_{\text{all}}<0$, indicating that the total resistance of the circuit has been optimized.
B.3 Analysis of Effectiveness of Chain-of-Self-consistency
Furthermore, we optimize the total circuit resistance from a different perspective, specifically by improving the resistance of validation resistor through self-consistency. In this approach, the validation resistor is decomposed into a parallel combination of multiple validation resistors and a meta-validation resistor with resistance $R_{\text{meta}}$, which verifies the correctness of the validation process. The original equation is thus reformulated as:
$R^{\text{CoV}}_{\text{all}}$ $=R_{0}+\frac{1}{\sum_{j=1}^{k}\sum_{i=1}^{n}\frac{1}{R_{\text{S}}}}+\sum_{j=1}^{k}\left(R_{\text{meta}}\right)+\frac{R_{\text{CoT}}}{n}$
$=R_{0}+\frac{R_{\text{S}}}{nk}+kR_{\text{meta}}+\frac{R_{\text{CoT}}}{n}.$
Clearly, as the sample size $n$ and the meta-validation steps $k$ increase, $\frac{R_{\text{S}}}{nk}$ decreases, leading to reduced difficulty in meta-validation. This significantly optimizes the total resistance and improves output power. Figure 6b illustrates this performance enhancement.
However, as the number of samples requiring validation grows, the model’s accuracy first increases and then decreases. Specifically, for sufficiently large validation steps, the resistance satisfies:
$\lim_{nk\to+\infty}R^{\text{CoV}}_{\text{all}}=R_{0}+kR_{\text{meta}}.$ (25)
With excessive validation steps, $k$ becomes too large, resulting in $kR_{\text{meta}}>R_{S}$. Consequently, $R^{\text{CoV}}_{\text{all}}$ exceeds its optimized value, leading to a performance trend characterized by initial improvement followed by decline.
C Proof of Impact of Resistor Optimization
To compute the derivatives of $P_{\text{model}}$ with respect to $R_{\text{CoT}}$, we start with the given expression:
$P_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{\text{CoT}}+R_{0})^{2}}.$ (26)
According to Eq. 10, let $R_{\text{CoT}}=R_{1}+R_{2}$, where $R_{1}\gg R_{2}$, if $R_{1}$ is improved to $\frac{R_{1}}{k}$, the power can be calculated as:
$P^{1}_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (27)
The corresponding power improvement is:
$\Delta P^{1}_{\text{model}}$ $=P^{1}_{\text{model}}-P_{\text{model}}$ (28)
$=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (29)
$=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}})$ (30)
Similarly, if $R_{2}$ is improved to the original $\frac{R_{2}}{k}$, the power improvement can be calculated as:
$\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (31)
To determine which improvement has a greater effect, consider:
$\Delta P^{1}_{\text{model}}-\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}(\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (32)
Define:
$A=\frac{R_{1}}{k}+R_{2}+R_{0},\quad B=R_{1}+\frac{R_{2}}{k}+R_{0}$ (33)
The fractional difference simplifies as:
$\frac{1}{A^{2}}-\frac{1}{B^{2}}=\frac{B^{2}-A^{2}}{A^{2}B^{2}}$ (34)
We need to prove that for $R_{1}\gg R_{2}$ whether $B^{2}-A^{2}$ larger or smaller than zero. Now, let’s calculate $B^{2}-A^{2}$:
$B^{2}-A^{2}=(B-A)(B+A)$ (35)
where:
$B-A$ $=\left(R_{1}+\frac{R_{2}}{k}+R_{0}\right)-\left(\frac{R_{1}}{k}+R_{2}+R_{0}\right)$ (36)
$=R_{1}-\frac{R_{1}}{k}+\frac{R_{2}}{k}-R_{2}$ (37)
$=R_{1}\left(1-\frac{1}{k}\right)+R_{2}\left(\frac{1}{k}-1\right)$ (38)
and:
$B+A$ $=\left(R_{1}+\frac{R_{2}}{k}+R_{0}\right)+\left(\frac{R_{1}}{k}+R_{2}+R_{0}\right)$ (39)
$=R_{1}+\frac{R_{1}}{k}+R_{2}+\frac{R_{2}}{k}+2R_{0}$ (40)
Analysis of the Sign of $B^{2}-A^{2}$:
$B^{2}-A^{2}=(B-A)(B+A)$ (41)
We only need to examine the sign of $B-A$, since $B+A>0$ holds true at all times.since $R_{1}\gg R_{2}$, and $01-\frac{1}{k}>$ 0 for $n>1$, we have:
$R_{1}\left(1-\frac{1}{k}\right)-R_{2}\left(1-\frac{1}{k}\right)$ $>0$ (42)
$(R_{1}-R_{2})\left(1-\frac{1}{k}\right)$ $>0$ (43)
$B^{2}-A^{2}$ $>0$ (44)
$\Delta P^{1}_{\text{model}}-\Delta P^{2}_{\text{model}}$ $>0$ (45)
Thus, $\Delta P^{1}_{\text{model}}<\Delta P^{2}_{\text{model}}$ the difference is negative, and the proof is complete. Therefore, The greater the original difficulty of the task corresponding to the capability, the greater the impact on the final performance.

### v2

@@
-ECM: A Unified Electronic Circuit Model for Explaining the Emergence of In-Context Learning and Chain-of-Thought in Large Language Model
+Electronic Circuit Principles of Large Language Models
 Abstract
-Recent advancements in large language models (LLMs) have led to significant successes across various applications, where the most noticeable is to a series of emerging capabilities, particularly in the areas of In-Context Learning (ICL) and Chain-of-Thought (CoT).
-To better understand and control model performance, many studies have begun investigating the underlying causes of these phenomena and their impact on task outcomes.
-However, existing explanatory frameworks predominantly focus on isolating and explaining ICL and CoT independently, leading to an incomplete understanding of their combined influence on model performance.
-To address this gap, we propose the Electronic Circuit Model (ECM), which provides a foundation for developing scalable, learnable policies and improving the management of AI-generated content.
-Specifically, ECM conceptualizes model behavior as an electronic circuit: ICL is represented as semantic magnetic field to providing an additional voltage following Faraday’s Law, while CoT is modeled as series resistors to constrain the model output performance following Ohm’s Law.
-Experimental results demonstrate that the ECM effectively predicts and explains LLM performance across a variety of prompting strategies.
-Furthermore, we apply ECM to advanced reasoning strategy optimization on a series of tasks, such as the International Olympiad in Informatics (IOI) and the International Mathematical Olympiad (IMO), achieving competitive performance that surpasses nearly 80% of top human competitors.
-Key Words: Electronic Circuit Model, Large Language Model, In-Context Learning, Chain-of-Thought
+Large language models (LLMs) such as DeepSeek-R1 have achieved remarkable performance across diverse reasoning tasks.
+To uncover the principles that govern their behaviour, we introduce the Electronic Circuit Principles (ECP), which maps inference-time learning (ITL) onto a semantic electromotive force and inference-time reasoning (ITR) onto a resistive network governed by Ohm’s and Faraday’s laws.
+This circuit-based modelling yields closed-form predictions of task performance and reveals how modular prompt components interact to shape accuracy.
+We validated ECP on 70,000 samples spanning 350 reasoning tasks and 9 advanced LLMs, observing a about 60% improvement in Pearson correlation relative to the conventional inference-time scaling law.
+Moreover, ECP explains the efficacy of 15 established prompting strategies and directs the development of new modular interventions that exceed the median score of the top 80% of participants in both the International Olympiad in Informatics and the International Mathematical Olympiad.
+By grounding LLM reasoning in electronic-circuit principles, ECP provides a rigorous framework for predicting performance and optimising modular components.
+Key Words: Electronic Circuit Principles, Large Language Model, Inference-time Learning, Inference-time Reasoning
+Figure 1: The Introduction of Inference-Time Learning (ITL) and Inference-Time Reasoning (ITR), and Comparison of Micro- and Macro-Circuits.
+a, Introduction to ITL, which pairs queries with solution in input context as references, and to ITR, which provides step-by-step reasoning in model-generated output.
+b, Electronic Circuit Principles on Macro-circuit analyses: Traditional micro-circuits capture basical linguistic functions but miss higher-order reasoning principles.
+The micro-circuit structure follows Rai et al.
+[35].
 1 Introduction
-Large Language Models (LLMs) have made significant strides across various tasks, including code generation, knowledge-based question answering, and complex reasoning [46, 3, 29, 32].
-Unlike earlier models that showed incremental improvements, LLMs exhibit emergent capabilities that fundamentally reshape problem-solving.
-At the core of these capabilities are two critical mechanisms: In-Context Learning (ICL) [2, 11, 31] and chain-of-Thought (CoT) reasoning [26, 41, 5], enabling LLMs to adapt dynamically to new challenges and achieve human-comparable learning and reasoning [2, 27, 1].
-Specifically, ICL triggers LLMs to generalize information based on task-specific context provided in the input, without requiring parameter updates.
-In parallel, CoT reasoning allows LLMs to approach complex problems stepwise, enhancing both performance and interpretability.
-Together, these mechanisms empower LLMs to serve as versatile problem solvers in critical fields such as healthcare and education [14, 25, 16, 18].
-Despite their transformative potential, the mechanism of LLMs’ emergent capabilities remains poorly understood.
-Current research on ICL has uncovered preliminary evidence of how models encode and process contextual information, with studies suggesting latent representations that approximate forms of task-specific meta-learning [38, 36] or topic models [39].
-Similarly, CoT research has begun to map the reasoning pathways LLMs employ, providing insights into their demonstration alignment with human cognitive heuristics [24, 37].
-Others seek to interpret emergent behaviors using macroscopic mathematical models based on experimental analysis [13, 5].
-In practice, ICL and CoT are often integrated, allowing models to adapt to context and solve problems step-by-step [17, 15, 30, 47, 6].
-To put it figuratively, combining ICL and CoT is like planting a tree: ICL provides the rich “soil” in which the model can absorb and integrate contextual knowledge, while CoT guides its “branches” to grow step by step in a clear, structured direction.
-Together, they create a natural, human-like reasoning process that’s both flexible and transparent, enabling models to tackle complex tasks with greater depth and precision.
-Unfortunately, in the current literature, a unified understanding of the combined effect of ICL and CoT remains limited, which seriously affects correct performance interpretation and controllable result prediction and further constrains the development of principled methodologies for LLM design and deployment.
-Figure 1: Glossary of Terms and Schematic of Electronic Circuit Model (ECM).
-In-Context Learning (ICL) is conceptualized as a (Semantic) Magnetic Field.
-Demonstrations in ICL effectively provide an additional voltage by applying a magnetic field change aligned with the input query direction.
-Chain-of-Thought (CoT) process is analogized as series resistors in circuit, where the reasoning difficulty in CoT are divided into a sequence of sub-difficulties, analogous to resistors in a circuit that collectively increase overall resistor.
-Building on this, we introduce the concept of model inherent capabilities, represented as the power supply of the circuit to provide basic voltage.
-Model performance can then be quantified as the circuit’s output power, enabling more precise predictions of empirical accuracy.
-In addition, in this article, “strength” denotes the value of semantic magnetic fields, “resistance” denotes the value of resistor, and “voltage” denotes the value of power supply.
-To bridge this gap, we introduce the Electronic Circuit model (ECM), a unified theoretical paradigm that reimagines model performance through the lens of physical circuit systems.
-Specifically, as shown in Figure 1, we try to conceptualize ICL capabilities as semantic magnetic fields, driving additional voltage gains that correspond to enhanced information absorption, while CoT challenges are modeled as series resistors, introducing predictable constraints akin to circuit losses.
-Based on these hypotheses, we discover Faraday’s Law of ICL and Ohm’s Law of CoT in LLMs’ reasoning process.
-By rigorously aligning theoretical circuit power predictions with empirical accuracy results across 3 prompting strategies, 4 task categories, and 13 different LLMs.
-The ECM successfully establishes a unified quantitative paradigm for predicting and optimizing the behavior of LLMs.
-Moreover, based on the introduction of ECM, we further explain over 10 existing prompt optimization strategies from the perspective of optimizing the magnetic field and optimizing the resistors.
-And given these perspectives, some phenomena that have not been successfully explained by previous works and more excellent prompt strategies are discovered based on ECM.
-Notably, our results show that combined with these approaches has achieved breakthroughs in complex domains: LLMs guided by ECM-optimized strategies have outperformed over 80% of human competitors in the International Olympiad in Informatics and the International Mathematical Olympiad.
-Furthermore, in subjective and exploratory contexts such as academic research development, ECM-optimized strategies have driven significant gains, achieving a minimum 10% improvement in overall human scores.
+Large Language Models (LLMs), like DeepSeek-R1 [14] and OpenAI-o1 [17], have recently achieved unprecedented performance in code generation, knowledge-based question answering and complex reasoning tasks [6, 25, 34].
+Unlike conventional software, whose behavior is explicitly encoded by human, LLMs exhibit emergent, stochastic dynamics that defy straightforward performance modeling.
+As a result, users often cannot anticipate when a model will succeed or fail before inference and generation, creating a “Schrödinger’s cat” scenario in which reliability only becomes apparent post hoc.
+At inference, as shown in Figure 1a, two complementary behaviors drive LLM adaptability [6]: inference-time learning (ITL) [2, 11, 33] and Inference-Time Reasoning (ITR) [30, 43, 4].
+ITL, often termed in-context learning, enables parameter‐free generalization from examples supplied in the prompt.
+By contrast, ITR, frequently implemented via chain-of-thought prompting, decomposes complex tasks into interpretable reasoning steps, enhancing both accuracy and transparency.
+These paradigms have underpinned rapid advances in domains ranging from clinical decision support to personalized education [15, 29, 20, 22], yet they also obscure the behind rationales governing performance, impeding systematic understanding and theory-guided refinement of reasoning strategies from higher-order perspectives.
+To address this gap, we present the Electronic Circuit Principle (ECP), which maps ITL and ITR onto analogues of classical circuit elements.
+As illustrated in Fig.
+1b, in this principle, ITL acts as an inference-time‐varying electromotive force, enabling demonstration incorporation much like induced voltage.
+Meanwhile, ITR maps onto an electrical network in which each reasoning subtask contributes series resistance, such that the total resistance quantifies cumulative cognitive load.
+By formalizing Faraday’s and Ohm’s laws of circuits, ECP furnishes a principled framework that both predicts LLM performance and guides modular optimization of reasoning components.
+We validate this approach on nine widely used LLMs across 350 tasks (over 70,000 samples), achieving a Spearman correlation > 0.7 between ECP‐based predictions and empirical accuracy.
+This represents about 60% correlation improvement over current inference-time scaling law methods that account for post-inference results, highlighting the enhanced predictive capacity of ECP.
+Moreover, ECP‐inspired adjustments, treating demonstration strength optimization as “voltage” improvement and reasoning strategy optimization as “resistance” reduction, consistently improve model performance based on theoretical insights.
+Notably, LLMs employing ECP-optimized strategies surpassed 80% of human competitors in the International Olympiad in Informatics (IOI) and the International Mathematical Olympiad (IMO) based on much weaker LLMs.
+Furthermore, in subjective and exploratory domains such as academic research, ECP-optimized strategies contributed to a minimum 10% improvement in human performance metrics.
+These findings demonstrate the utility of the Electronic Circuit principle for targeted modular optimization.
 In summary, our contributions are as follows:
-  • To the best of our knowledge, we are taking the first meaningful step to introduce the Electronic Circuit Model (ECM), effectively offering a unified theoretical paradigm to reformulate model performance through the lens of circuit systems.
-  • In this framework, we discover Faraday’s Law of ICL and Ohm’s Law of CoT in LLMs’ reasoning, and further utilize ECM to effectively predict the empirical accuracy and explain the effectiveness of 10 existing strategies from the view of optimizing the magnetic field and optimizing the resistor.
-  • We utilize ECM to explain some open phenomena and to propose more excellent prompt strategies based on ECM.
-Combined with these approaches, we have achieved breakthroughs in mathematical and code competition and academic research development scenarios.
-Further, we empirically demonstrate that applying ECM-driven optimizations to LLMs leads to substantial performance gains.
-2 Theoretical Model
-This theoretical model shows that model performance ($P_{\text{out}}$) relies not only on basic voltage from model inherent capabilities ($\mathcal{E}_{\text{model}}$) but also on external factors like reasoning difficulty ($R_{CoT}$) and additional voltage from extra demonstrations in ICL ($\mathcal{E}_{ICL}$), which are described as follows in detail:
+  • Uncovering Macro-level Principles in LLMs: We introduce the Electronic Circuit Principle (ECP), a pioneering framework that redefines model performance through the lens of macro-circuit, providing a unified theoretical paradigm for quantitatively predicting and optimizing LLM performance.
+  • Pioneering Accurate Performance Prediction: By drawing analogies to Faraday’s Law of Induction and Ohm’s Law, we offer new insights into LLM reasoning mechanisms, enabling precise performance predictions with a Spearman correlation exceeding 70% from 9 widely-used LLMs across more than 350 tasks.
+  • Guiding Modular Performance Optimization: Leveraging ECP, we interpret existing optimization strategies and propose novel optimization strategies informed by ECP theory.
+These strategies have led to significant breakthroughs in various fields, from mathematical and programming competitions to cutting-edge academic research.
+Figure 2: Glossary of Terms and Schematic of Electronic Circuit Principles (ECP).
+Inference-time learning (ITL) is depicted as a dynamic semantic magnetic field whose strength rises when demonstrations accompany the prompt, adding potential that steers token selection.
+Inference-time reasoning (ITR) is modelled as a series circuit: each reasoning sub-task acts as a resistor, and their sum gives the total cognitive load.
+The LLMs’ intrinsic capacity supplies the driving voltage, and empirical accuracy corresponds to the power delivered across the load.
+“Strength”, “Resistance” and “Voltage” map to field intensity, sub-task difficulty and LLM capability, respectively, offering a compact framework to quantify and compare LLM behaviour.
+2 Electronic Circuit Principles
+Electronic Circuit Principles show that model performance ($P_{\text{out}}$) relies not only on basic voltage from model inherent capabilities ($\mathcal{E}_{\text{model}}$) but also on external factors like reasoning difficulty ($R_{ITR}$) and additional voltage from extra demonstrations in ITL ($\mathcal{E}_{ITL}$), which are described as follows in detail:
 2.1 Model Inherent Capabilities as Power Supply
-In our model, we draw an analogy by conceptualizing LLM’s inherent capabilities as a power supply that drives computational processes and reasoning tasks (Figure 1).
+In ECP, we draw an analogy by conceptualizing LLM’s inherent capabilities as a power supply that drives computational processes and reasoning tasks (see in Fig.
+2).
 The voltage value $\mathcal{E}_{\text{model}}$ represents the strength of the inherent capability, governing the LLM’s capacity to execute a range of tasks.
 Just as the voltage in an electrical circuit determines the power available to drive various components, the internal capabilities govern its ability to manage increasingly complex inputs.
-2.2 Faraday’s Law of In-Context Learning
+2.2 Faraday’s Law of Inference-Time Learning
 Inspired by Wang et al.
-[38], the process of ICL can be conceptualized as the interaction of LLM with contextual data, where it absorbs, retains, and eventually releases semantic information flow.
-This dynamic flow of information resembles the changing behavior of a magnetic field as its initial strength $\Phi^{B}_{0}$ decreases to zero.
-Building on this analogy, we propose Faraday’s Law for the ICL paradigm, where the rate of change in the “magnetic field" further induces an extra voltage:
-$\mathcal{E}_{\text{ICL}}=-\frac{d\Phi^{B}(t)}{dt}=\lambda\Phi_{0}^{B},$ (1)
-where $\lambda$ represents the uniform rate of decrease in magnetic field intensity.
-This perspective enables a deeper quantitative understanding of ICL through the concept of a “semantic magnetic field," as illustrated in Figure 1.
-2.3 Ohm’s Law of Chain-of-Thought
-In a parallel vein, Chen et al.
-[5] have discovered a combination law for LLM reasoning, which aligns with the series formula of electrical resistors after transformation (See in Appendix A).
-Inspired by this, we conceptualize the difficulty of CoT reasoning as analogous to series circuits.
-In this framework, each reasoning process introduces a distinct resistor ($R_{\text{CoT}}=\sum_{i}R_{i}$), compounding the task’s overall difficulty ($R_{i}$).
-As illustrated in the right panel of Figure 1, Ohm’s law for CoT can be expressed as:
-$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}}{R_{\text{CoT}}+R_{0}},$ (2)
+[40], the process of ITL can be conceptualized as the interaction of LLM and contextual data, whereby the model absorbs, retains, and ultimately releases semantic information.
+This dynamic flow of information is analogous to the behavior of a magnetic field, where its initial strength, $\Phi^{B}_{0}$, diminishes to zero over time.
+Based on this analogy, as illustrated in Fig.
+2 (left), we introduce Faraday’s Law within the context of ITL, where the rate of change in the "magnetic field" induces a corresponding voltage.
+Specifically, we define the voltage as:
+$\mathcal{E}_{\text{ITL}}=-\frac{d\Phi^{B}(t)}{dt}=\lambda\Phi_{0}^{B},$ (1)
+where $\lambda$ denotes the uniform rate of magnetic field intensity decay. This framework provides a quantitative approach to understanding ITL by the concept of a “semantic magnetic field”.
+2.3 Ohm’s Law of Inference-Time Reasoning
+Similarly, Chen et al.
+[4] have derived a combination law for reasoning within LLM, which parallels the series formula for electrical resistors (See Methods).
+Drawing on this, we propose a principle for the difficulty of ITR that is analogous to series circuits.
+In this principle, each reasoning process introduces a distinct resistor ($R_{\text{ITR}}=\sum_{i}R_{i}$), compounding the task’s difficulty ($R_{i}$).
+As illustrated in Fig.
+2 (right), Ohm’s law for ITR follows:
+$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}}{R_{\text{ITR}}+R_{0}},$ (2)
 where $R_{0}$ represents the resistance of an output resistor with static resistance value in the circuit, which reflects the difficulty of reasoning conclusion.
-This formulation quantifies the cumulative effect of reasoning resistors, providing a modular quantitative perspective on cognitive task complexity.
+This formulation quantifies the cumulative effect of reasoning resistors, providing a modular quantitative perspective on task complexity.
 2.4 Model Performance as Output Power of Bulb
-Extending the conceptual model, we equate model performance to the power output of a light bulb in an electronic circuit.
+Extending the ECP, we equate model performance to the power output of a light bulb in an electronic circuit.
 Task execution efficiency depends on a nonlinear interaction of internal and external factors, rather than solely on the model’s intrinsic capabilities.
-The output power, $P_{\text{out}}$, is expressed as:
-$P_{\text{out}}=I^{2}_{model}R_{0}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{\text{CoT}}+R_{0})^{2}}.$ (3)
+The output power, $P_{\text{out}}$, can be formally expressed as:
+$P_{\text{out}}=I^{2}_{model}R_{0}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{\text{ITR}}+R_{0})^{2}}.$ (3)
 Here, output power represents the model’s effective performance, with higher power indicating better task execution, akin to a brighter bulb.
-The equation highlights that maximizing performance requires enhancing the basic voltage of power supply ($\mathcal{E}_{\text{model}}$) of LLM, minimizing reasoning resistor ($R_{\text{CoT}}$) in CoT, and optimizing the extra voltage ($\mathcal{E}_{\text{ICL}}$) provided by ICL.
+The equation highlights that maximizing performance requires enhancing the basic voltage of power supply ($\mathcal{E}_{\text{model}}$) of LLM, minimizing reasoning resistor ($R_{\text{ITR}}$) in ITR, and optimizing the extra voltage ($\mathcal{E}_{\text{ITL}}$) provided by ITL.
 This synergy explains and predicts the model’s performance on complex tasks quantitatively.
-Figure 2: The Experimental Results of Theoretical Model Verification on ECM.
-a,
-An analysis of ECM incorporating the CoT approach reveals a scatter plot illustrating the relationship between model accuracy and theoretical power with 0.8843 Spearman coefficient.
-b.
-Six similarity metrics are employed to evaluate whether context demonstrations can be interpreted as a semantic magnetic field.
-The results indicate that the projection length achieving the highest coefficient of 0.9106 between practical accuracy and theoretical power.
-c, The integrated ECM framework combines ICL and CoT.
-Across three retrieval strategies for ICL demonstrations, a consistent linear correlation is observed.
-d, Experiments on various LLMs under different power supply voltages, representing their capacities, reveal a strong linear correlation between theoretical power estimates and observed accuracy.
-e, Broader analyses across diverse tasks, including multi-hop knowledge reasoning, medical probing, and robotic planning, demonstrated robustness.
-3 Theoretical Model Verification
-3.1 CoT Meets Ohm’s Law
-First, we verify whether CoT adheres to Ohm’s law.
-Since the electric current $I_{model}$ in Eq.
-2 is not directly observable, we instead examine whether $P_{out}$ is linearly related to true accuracy, thereby indirectly validating CoT Ohm’s law.
-Specifically, to control variables, we remove the ICL and let the model perform zero-shot reasoning step-by-step following the instructions, which satisfies: $P_{\text{out}}=\frac{\mathcal{E}_{\text{model}}^{2}R_{0}}{(R_{\text{CoT}}+R_{0})^{2}}.$ The strong Spearman correlation coefficient between theoretical power ($P_{\text{out}}$) and observed accuracy (Acc), shown in Figure 2a, is 0.8843, with a $p$-value below 0.01.
-This strong correlation suggests difficulty of CoT reasoning actually exhibits resistor-like behavior, making it a reliable predictor of model performance.
-3.2 ICL Meets Faraday’s Law
-We conceptualize contextual demonstrations in ICL as a semantic magnetic field governed by Faraday’s Law during reasoning.
-Under this framework, semantic influence and proximity are quantified by projection length, as illustrated in the lower right of Figure 2b.
-To validate this hypothesis, we evaluate various metrics for contextual proximity between demonstrations and user requests.
-Our analysis (Figure 2b) shows that the Spearman correlation between the calculated power $P_{\text{model}}$ in Eq.
-3 and observed spearman coefficient and R2 merely for projection length exceeds 0.85.
-This indicates that the geometric alignment of the demonstrations with the task request is a strong predictor of model performance, which further prove the existance of Faraday’s Law of ICL.
-3.3 ICL and CoT Satisfy a Unified ECM
-In an unprecedented integration, we unify ICL and CoT within a single ECM, offering a comprehensive framework to explain their interaction and collective impact on model performance.
-We rigorously test this unified model using three distinct retrieval strategies for ICL demonstrations, all of which yields a consistent linear distribution, as illustrated in Figure 2c.
-This consistent alignment, with a Spearman correlation coefficient of 0.8979 and a $p$-value approaching zero between calculated power in Eq.
-3 and practical accuracy, provides strong empirical evidence for the robustness of our theoretical model.
-These results confirm that both ICL and CoT not only can be calculated independently but are inherently compatible within the same theoretical structure, offering a powerful new way to describe and predict the behavior of LLM systems.
-3.4 Extension Experiments on Various LLMs and Tasks.
-To demonstrate the generality of our theoretical ECM, we conduct extension experiments across LLMs configured with varying power supply voltages to represent different capacities.
-Figure 2d reveals a consistent linear correlation over 0.8348 between theoretical output power estimates and empirical accuracy, confirming the ECM’s robustness in predicting performance.
-Beyond traditional mathematical reasoning tasks, we test ECM on diverse real-world problems, including multi-hop knowledge reasoning [44], medical probing [7], and robotic planning [34].
-As shown in Figure 2e, ECM achieves Spearman coefficients exceeding 0.86 and R2 exceeding 0.77 across all tasks, underscoring its versatility in both cognitive and practical applications.
-4 Explanation based on ECM
-In this section, we mainly pay attention to how to use ECM to explain the effectiveness of existing prompting methods.
-4.1 Explanation for ICL Strategies
-Enhanced embeddings facilitate accurate semantic magnetic field strength estimation and retrieval.
-Recent advancements in embedding models have shown that improved text representations directly enhance LLM’s efficacy in the retrieval process [21].
-This improvement is tied to the precise estimation of semantic magnetic field strength.
-As shown in Figure 3a, we employ three models—BERT [9], RoBERTa [22], and BGE [4]—to represent and retrieve demonstrations in ICL.
-Among these, BGE achieves the highest degree of alignment, with a Spearman correlation coefficient of 0.9106 between theoretical power and empirical accuracy, significantly surpassing the other models.
-This result highlights superior embedding ability attributed to the accurate semantic magnetic field estimation.
-Figure 3: The Experimental Explanation for Existing ICL Strategies.
-a,
-The relationship between theoretical power and actual accuracy for various encoders is assessed.
-BGE achieves the highest Spearman correlation coefficient of 0.9106.
-b, A comparison of the Vanilla Encoder and the Meta-Recognition Representation reveals the latter’s ability to enhance the linear correlation between theoretical power and practical accuracy, from 0.9130 to 0.9337.
-c, The distribution of magnetic field strengths in generated demonstrations (with CDF representing cumulative distribution function and CCDF representing complementary cumulative distribution function).
-Further, we demonstrates the effectiveness of different Synthetic-CoT with degree of different degree of semantic field strength.
-d, The effectiveness of Synthetic-CoT is analyzed across three conditions: Positive Synthetic-CoT (demonstrations with positive extra voltage from ICL based on positive magnetic fields), Zero-shot CoT (no demonstrations), and Negative Synthetic-CoT (demonstrations with negative extra voltage from ICL based on negative magnetic fields).
-The left scatter plot displays the performance distribution of these conditions on the power-accuracy landscape.
-The right bar chart highlights factors contributing to Synthetic-CoT’s optimization of extra voltage provided by ICL.
-Meta Recognition Enhance Magnetic Field Strength Estimation Modeling.
-The study of Didolkar et al.
-[10] introduces Meta-Recognition, a method leveraging GPT-4 [1] to decompose complex problems into modular sub-capabilities for ICL representation, improving LLMs’ performance.
-We attribute its effectiveness to the enhanced sample representation modeling, which enhances ICL voltage and subsequently boosts output power and accuracy.
-As illustrated in Figure 3b, Meta-Recognition significantly increases the correlation between theoretical power and practical accuracy, improving from 0.9130 to 9337.
-It underscores that representation based on sub-capability decomposition effectively enhances the precision of magnetic field strength modeling.
-Synthetic-CoT predominantly generates samples with positive magnetic fields, enhancing model performance.
-Synthetic-CoT, a widely used approach in ICL, enables LLMs to autonomously generate ICL demonstrations, thereby improving performance [33].
-As illustrated in Figure 3d, all ICL demonstrations produced by Synthetic-CoT exhibit positive semantic magnetic field contributions, which are strongly correlated with enhanced model accuracy.
-This effect is likely due to the positive voltage induced by the magnetic field, which optimizes circuit functionality.
-As shown in Figure 3d, the larger ICL magnetic field strength corresponds to the larger power and the higher accuracy.
-To further investigate the impact of magnetic field polarity on model performance, we analyzed demonstrations with negative magnetic field.
-Including such examples led to a 21.47% reduction in accuracy and a substantial decline in theoretical power output (Figure 3d).
-These findings underscore the critical role of positive magnetic field demonstrations in the superior performance achieved by Synthetic-CoT.
-Figure 4: The Experimental Explanation for Existing CoT Strategies.
-a,
-The distributions of accuracy, power, and resistance for Zero-shot without CoT and Zero-shot with CoT are presented.
-The average resistance for Zero-shot without CoT is 10.93, significantly higher than the 1.50 observed for Zero-shot with CoT.
-b, Mechanistic analysis of Tool Usage (TU) and Program-of-Thought (PoT) within the ECM.
-Initially, TU, modeled on the original CoT, had non-zero computational resistance ($R_{C}$), limiting power-accuracy correlation.
-Assuming $R_{C}=0$ achieves a 10% improvement in the Spearman coefficient.
-PoT, representing TU with structured code, further enhanced performance by keeping $R_{C}=0$ and reducing planning resistance $R_{P}$, achieving another 10% improvement in power-accuracy correlation.
-c, A circuit-based rationale to explain how to utilize self-consistency and coverage to enhance performance.
-Self-consistency reduces total resistance by processing multiple reasoning resistors in parallel (sampling $n$ solutions and selecting the output with the highest consensus).
-In contrast, coverage minimizes resistance by increasing sample size with negligible resistance of voting resistor (sampling $n$ instances and passing the task if any succeeds).
-4.2 Explanation for CoT Strategies
-Direct Answering Introduces Greater Logical Resistor Compared to CoT
-Recent studies suggest that generating direct answers without intermediary reasoning steps increases cognitive load on the model [13, 17].
-We hypothesize that the absence of explicit reasoning pathways introduces significant resistance of “logical resistor,” leading to weaker performance compared to Zero-shot CoT.
-To test this, we analyzed performance and power correlations for direct answering and Zero-shot CoT on a power-accuracy graph, finding a strong linear correlation of 0.9131 ($p$-value $<0.01$).
-The average resistance for direct answering is calculated as 10.93, substantially higher than the 1.50 observed for Zero-shot CoT.
-This discrepancy corresponds to a performance reduction of over 30% for direct answering.
-These findings indicate that the lack of a CoT framework imposes additional greater logical resistor, hindering the model’s reasoning capabilities and resulting in diminished performance.
-Tool Usage optimizes computational resistor, and PoT further reduces planning resistor, improving model performance.
-Chen et al.
-[5] propose that Program of Thought (PoT) and Tool Usage (TU) enhance computational reasoning capabilities to a theoretical maximum, effectively minimizing reasoning complexity to zero, akin to eliminating the resistance to zero.
-Inspired by this, we investigate the correlation between CoT and TU models under this "zero-resistance" assumption.
-Our analysis revealed that modeling TU as a circuit with zero resistance of computational resistor increased the Power-accuracy Spearman correlation by over 10% compared to the original model, as illustrated in Figure 4b, which validates our hypothesis.
-Furthermore, we hypothesize that PoT, leveraging programming languages, reduces planning resistance.
-As shown in Figure 4b, modeling PoT as a circuit with lower planning resistance achieved a correlation score exceeding 0.9423 between theoretical power and accuracy—significantly higher than TU.
-This finding highlights the superior effectiveness of PoT in improving model performance.
-Self-consistency and Inference Scaling Law via Parallel Resistor in Circuits
-Self-consistency, as demonstrated by Wang et al.
-[40], utilizes multiple reasoning paths simultaneously, and determines final decision through majority voting, enhancing performance robustness and accuracy.
-Here, we interpret its effectiveness using a circuit model, which introduce parallelization of original resistors ($R_{\text{CoT}}$), significantly reducing the total system resistance.
-As shown in Figure 4c, the practical accuracy of the self-consistency strategy aligns closely with the theoretical output power curve, validating the correctness of our proposed model.
-Recent findings on the Inference Scaling Law [43, 28] further demonstrate that increasing the number of parallel samples, $n$, substantially improves accuracy.
-As $n$ grows, the resistance value of CoT resistor approaches zero, optimizing the reasoning process.
-Further, when coverage introduces an absolutely correct answer verification process with zero resistance of verification resistor, exceeding the gains achieved by self-consistency alone.
-The alignment between theoretical predictions and empirical results, also evident in Figure 4c, reinforces the validity of this model.
-These results highlight the central role of self-consistency and inference scaling in enhancing the performance of circuits.
-Figure 5: The Experimental Exploration for ICL Strategies.
-a,
-The relationship between the introduction of demonstrations and model performance.
-Using Reverse ICL + CoT, which introduces demonstrations with negative magnetic fields, impairs the model’s reasoning ability and results in performance falling below that of Zero-shot CoT.
-In contrast, Retrieval-ICL + CoT, incorporating demonstrations with positive magnetic fields, significantly improves performance.
-b, The upper part of the figure contrasts Diversity-oriented Static ICL (DS-ICL), where demonstrations are retrieved for the whole task, and Similarity-oriented Dynamic ICL (SD-ICL), where demonstrations are retrieved for each request.
-In SD-ICL, demonstrations are more aligned with test requests, creating a stronger semantic magnetic field.
-The lower part highlights Diversity-oriented Dynamic ICL (DD-ICL), where increasing sampling range enhances the diversity but decreases the magnetic field strength of retrieved demonstrations, resulting in a more diverse but weaker strength, leading to less accuracy.
-c, The left figure illustrates the demo-optimization process, a methodology designed to refine the demonstrations used in ICL.
-The right figure shows that demo-optimization outperforms traditional prompt optimization by improving theoretical power.
-5 Exploration based on ECM
-In this section, we primarily focus on how to use ECM to explore phenomena that have never been well explained and more novel prompting methods.
-5.1 ICL Exploration
-When Few-shot CoT Performs Worse than Zero-shot CoT?
-Despite the growing importance of ICL and CoT, a critical question remains: under what conditions is ICL essential, and when does zero-shot CoT suffice?
-To investigate this, we systematically examine the interaction between these two strategies.
-As shown in Figure 5a, our findings reveal a phenomenon we term “reverse ICL," where demonstrations associated with a negative semantic magnetic field significantly impair the output power of circuits, thereby decreasing the model’s reasoning ability and resulting in worse performance compared to zero-shot CoT.
-This suggests that not all demonstrations contribute positively to the learning process.
-In contrast, demonstrations with a positive semantic magnetic field lead to an enhancement in performance.
-These results highlight the critical importance of selecting demonstrations based on their associated semantic magnetic field, which plays a crucial role in determining the success of ICL.
-The influence of diversity in ICL retrieval: a static vs. dynamic perspective
-In ICL retrieval, two paradigms are commonly studied: static ICL, where the set of demonstrations remains fixed during testing, and dynamic ICL, where demonstrations are selected in real-time based on the specific test sample.
-Previous studies [19] reveal that semantic diversity in static ICL enhances performance by exposing the model to a broader range of contextual cues.
-Conversely, in dynamic ICL, diversity does not provide the same benefits and may even hinder performance [31].
-This raises the question: Why does diversity yield such contrasting outcomes in these retrieval modes?
-To this end, we hypothesize that the benefits of diversity in static ICL can be likened to a magnetic field.
-Diverse demonstrations create a robust and expansive field, offering the model access to varied directional signals as test samples change.
-This broad-spectrum magnetic field stabilizes the model’s decision-making process, improving its generalizability and overall performance.
-As illustrated in Figure 5b, static ICL benefits from this diversity, which amplifies the model’s predictive power compared to random selection strategies.
-In dynamic ICL retrieval, however, the focus shifts toward selecting demonstrations closely aligned with the test sample.
-This approach optimizes the magnetic field for stronger, more targeted signals in the relevant direction.
-In this context, diversity becomes minimally critical; the alignment and intensity of retrieved demonstrations take precedence.
-As shown lower panel in Figure 5b, diverse but less aligned demonstrations enable less accuracy.
-Demo-Optimization: a novel approach to enhance demonstration efficacy in ICL
-Recent advancements in prompt-based learning leverage feedback from outputs as model gradients, enabling large models to autonomously refine their prompts for more efficient learning [45].
-Expanding on this foundation, we propose demo-optimization, a methodology tailored to optimize the demonstrations used in ICL.
-Demo-optimization focuses on balancing the relative effectiveness between user requests and demonstrations (See Method for implementation details).
-As illustrated in Figure 5c, demo-optimization significantly outperforms traditional prompt optimization approaches, which primarily refine instructions.
-This method not only improves accuracy but also theoretically strengthens reasoning capabilities.
-Moreover, it addresses the persistent underfitting problem in prior ICL methods, offering a more precise and scalable strategy for enhancing demonstration-driven learning.
-Figure 6: Experimental Evaluation of Chain-of-Thought Strategies and Power-Optimized Performance on Complex Tasks
-a,
-Overview of the fine-grained self-consistency strategy in circuits.
-By decomposing parallel CoT resistors into smaller units, this approach effectively reduces overall circuit resistance value.
-b, Overview of the Chain-of-Verfication (CoV) strategy in circuits.
-In addition to optimizing parallel CoT resistances, this method incorporates the concept of voting resistors in self-consistency to further reduce overall circuit resistance.
-The CoV strategy achieves this by sampling multiple voting results and validating them to finalize the decision.
-c, Leveraging stronger magnetic fields and reducing logical resistance significantly enhances circuit performance.
-The power-enhanced model achieved notable results in the IOI (left figure) and the IMO (right figure).
-d, The methodology for incorporating sufficient theoretical power into manual ratings during idea generation scenarios.
-This approach demonstrates a significant improvement over the original solution by enhancing the accuracy and reliability of the evaluation process.
-5.2 CoT Exploration
-Optimize the same proportion of capabilities. The greater the original difficulty of the task corresponding to the capability, the greater the impact on the final performance.
-We can prove that optimizing a resistor with higher resistance to its original value $\frac{1}{k}$ results in a greater reduction in overall resistance of reasoning resistor $R_{\text{CoT}}$, yielding a more significant performance improvement compared to optimizing a lower resistance (See in Theoretical Proof).
-As shown in the meta-analysis in Appendix A, when solving questions that demand advanced planning, o1 [28]’s reasoning method, which emphasizes extensive exploration, significantly enhances the most challenging reasoning components.
-This improvement ultimately boosts overall performance on complex reasoning tasks.
-However, for tasks where the hardest component requires extensive domain-specific knowledge, insufficient optimization of such knowledge limits performance gains in those domains.
-Fine-grained Self-consistency achieves better performance
+Figure 3: Verification of the ECP via zero‐shot reasoning.
+a, Log-linear relationship between the classical inference-time scaling law (log token length versus task accuracy) and empirical performance of 4 LLMs, GPT-3.5 Turbo, GPT-4o-mini, GPT-4o and Qwen2.5-7B-Instruct, reveals a significant negative correlation (Spearman $\rho\in[-0.5664,-0.4254]$,$P<0.01$).
+b, Observed zero-shot accuracy plotted against ECP-predicted computational power for 9 LLMs shows a strong positive correlation (Spearman $\rho>0.71$, $P<0.01$), confirming the ECP’s predictive validity.
+c, Comparative evaluation of 4 semantic-similarity metrics used to estimate field strength: projection length, Manhattan ($L_{1}$) distance, Euclidean ($L_{2}$) distance and cosine similarity, identifies projection length as the most accurate predictor.
+More complete results are shown in Fig.
+9 in Appendix.
+3 Results
+3.1 Accurate Performance Prediction
+3.1.1 ITR Meets Ohm’s Law to Accurately Predict Zero-shot Performance
+To investigate whether ITR obeys Ohm’s law, we eliminate all inference-time learning effects (i.e., we exclude any in-context demonstrations in the prompts), so that $\mathcal{E}_{\mathrm{ITL}}=0$.
+Since the model current $I_{\mathrm{model}}$ in Eq.
+2 be measured directly, we instead quantify the output power to assess ITR’s adherence to Ohm’s law, which can be calculated as follows:
+$Acc\propto P_{\mathrm{out}}=\frac{\mathcal{E}_{\mathrm{model}}^{2}\,R_{0}}{(R_{\mathrm{ITR}}+R_{0})^{2}}.$ (4)
+As shown in Fig.
+3a, the traditional Inference Time Scaling Law, predicting performance from token length alone, fails to predict reasoning performance accurately (Spearman’s $\rho\approx-0.5$ and Pearson’s $r<0.53$).
+In contrast, as illustrated in Fig.
+3b, the output power of our ECP correlates strongly and positively with actual accuracy across every model tested ($\rho,r>0.7$).
+Our findings reveal that reasoning difficulty functions analogously to electrical resistance: as cognitive “resistance” increases, the “current” of information flow, and the relevant “power” of reasoning output diminishes, echoing the dynamics of Ohm’s law.
+3.1.2 ITL Meets Faraday’s Law to Accurately Predict Few-shot Performance
+Next, we introduce a novel conceptual framework for ITL, drawing an analogy to Faraday’s law: contextual demonstrations can be interpreted as generating a “semantic magnetic field” that influences task performance.
+To probe this analogy, we compare four proximity metrics: projection length, a baseline with no distance modelling, Euclidean distance, and cosine distance, quantifying the relationship between demonstrations and task prompts.
+As shown in Fig.
+3c, projection length exhibits the strongest correlation with task accuracy ($\rho=0.8251$, $r=0.8481$, $R^{2}=0.6095$), outperforming both the No Distance baseline ($\rho=0.7913$, $r=0.8320$, $R^{2}=0.5555$) and conventional metrics such as Euclidean distance ($\rho=0.7597$, $r=0.8170$, $R^{2}=0.6017$) and cosine distance ($\rho=0.8151$, $r=0.8217$, $R^{2}=0.5190$).
+Strikingly, the latter two, even though widely used for demonstration retrieval, perform worse than having no distance modelling at all.
+These findings reveal that projection length between demonstrations and task requests is another key to effectively model ITL effectiveness.
+Figure 4: Interpreting for Inference‐Time Learning Strategies of ECP.
+a, ECP‐predicted theoretical power correlates positively with accuracy across three semantic encoders (BERT, RoBERTa, BGE) over 350 tasks.
+b, Meta‐recognition embeddings yield higher mean theoretical power (73.6 vs 66.9) and accuracy (54.8% vs 49.9%) than vanilla encoders, with stronger Spearman correlations (0.75 vs 0.61, $P<0.01$), which using RoBERTa as semantic encoders.
+c, Cumulative distributions (CCDF, CDF) and 95% confidence ellipses for three ITR configurations (Acc = 0.442, 0.376, 0.352) illustrate that larger induced semantic fields correspond to higher theoretical power and observed accuracy.
+3.2 Modular Performance Optimization
+3.2.1 Interpreting Existing Optimization Strategies by Magnetic Field Adjustment
+Enhanced semantic encoders sharpen magnetic-field estimates and thereby improve retrieval accuracy.
+Recent progress in neural semantic encoders has markedly improved the performance of ITL tasks [26].
+To understand the underlying reason, we compared three prominent encoders (BERT [9], RoBERTa [27], and the more recent BGE [3]), by correlating power-accuracy metrics (Fig.
+4a).
+BGE exhibited the strongest agreement with ground truth (Spearman’s $\rho=0.83$, Pearson’s $r=0.85$, $R^{2}=0.61$), surpassing its predecessors and indicating that higher-quality embeddings generate more coherent semantic fields and more reliable performance forecasts.
+Meta-recognition representations further augment magnetic-field modelling.
+[…]
+Synthetic-CoT consistently consistently produces positive magnetic fields.
+Synthetic-CoT requires LLMs to generate demonstrations at inference [37].
+Stronger fields correlate positively with retrieval accuracy, as shown by complementary cumulative distributions and 95% confidence ellipses in power–accuracy space (Fig.
+4c).
+High-field Synthetic-CoT achieved an accuracy of 0.442, outperforming its low-field counterpart (0.376) and the zero-shot baseline (0.352).
+These converging results identify field magnitude as a key determinant of ITL performance and outline scalable principles for tuning magnetic fields through representation learning and prompt engineering.
+Figure 5: Interpreting for Inference‐Time Learning Strategies of ECP.
+a, ITR reconfigures the model’s reasoning circuit: relative to Rationale, removing the rationale (w/o Rationale) increases effective resistance (3.97 $\Omega$ vs 4.40 $\Omega$), lowers theoretical power (62.1 vs 56.7) and improves accuracy (41.7% vs 33.6%), as shown by the contour shifts.
+b, Applying natural language (NL), tool usage (TU) or program-of-thought (PoT) to instantiate the “superconducting” phenomenon ($R_{c}=0$) surpasses NL for better Power-Accuracy correlations, in line with findings on externalized reasoning.
+c, Circuit analogies for two sampling strategies.
+Self-consistency sampling comprises parallel reasoning paths merging through an aggregation resistance $R_{s}$, denoting consensus cost; coverage sampling sets ($R_{s}=0$), allowing independent paths for greater diversity.
+d, Empirical trends follow a logarithmic relation with sample count.
+e, Adopting independent rather than mixture sample counts motivates replacing linear models with logarithmic ones, markedly improving predictions.
+3.2.2 Interpreting Existing Optimization Strategies through Resistor Reorganization
+Direct answering increases logical resistance relative to zero-shot ITR.
+Prompting LLMs to produce answers directly, without exposing their intermediate reasoning, imposes greater cognitive demands and reduces task performance [43, 21], which we attribute it to imposing additional “logical resistance” on the inference pathway.
+As illustrated in Fig.
+5a, we found that direct answering lowers the mean theoretical power requirement (62.10 $\rightarrow$ 56.65)but simultaneously raises effective logical resistance from 3.97 to 4.40, driven by 1.1×, 1.5× and 1.6× increases in planning, calculation and local‐operation resistances, respectively.
+Accuracy consequently declines from 41.7% to 33.6%, indicating that explicit chains of thought enhance energy efficiency and steer the model towards correct solutions.
+Tool usage and program-of-thought minimise calculation and planning loads.
+Following Chen et al.
+[4], representing reasoning complexity as an ohmic load, we tested whether explicit tool calls (tool usage, TU) and step-wise code generation (program-of-thought, PoT) approach a “superconducting” limit.
+As shown in Fig.
+5b, setting the computational resistor to zero improved the power–accuracy fit by $8\%$ in Pearson’s $r$ and by $10\%$, supporting the zero‐resistance “superconducting” hypothesis for TU.
+Further reducing the planning resistor for PoT produced an even tighter correspondence ($r>0.95$ and $R^{2}>0.89$), demonstrating PoT’s superior efficiency in converting computational throughput into correct answers.
+Self-consistency and inference scaling act as parallel resistors.
+As shown in Fig.
+5c, executing $n$ concurrent reasoning chains and selecting the majority answer, also known as Self-consistency [41], can be modelled as $n$ parallel resistors of value $R_{\text{ITR}}$ which reduce the effective resistance to $R_{\text{ITR}}/n$; a verification step of resistance $R_{S}$ sets a lower bound (See Theoretical Proof).
+Inference-time scaling-law analyses [45, 31] show that enlarging $n$ drives $R_{\text{ITR}}/n$ towards zero, asymptotically maximising accuracy, and that a perfect verifier ($R_{S}=0$) outperforms self-consistency alone.
+Empirically, however, Pass@$k$ diverges from theory if samples are assumed independent (Fig.
+5c (left)).
+Accounting for correlations reveals only $\log n$ independent trajectories; replacing $R_{0}/n$ with $R_{0}/\log n$ restores perfect agreement between prediction and experimental data (Fig.
+5c (right)), confirming that the equivalent-circuit framework captures the diminishing returns of modern inference-time scaling-laws.
+Figure 6: Optimization Exploration through Magnetic Field Adjustment on ECP.
+a, Comparison of accuracy and theoretical power for no‐demo, bottom‐$k$ and top‐$k$ settings, top‐$k$ retrieves for the largest extra voltage.
+b, Retrieval strategy comparison for random, Diversity-oriented static ICL (DS-ICL) and Similarity-oriented dynamic ICL (SD-ICL) illustrate that SD-ICL achieves the greatest accuracy and power; line plots show both metrics decline as $k$ increases based on Diversity-oriented dynamic ICL (DD-ICL).
+c, Compared to manual ICL, optimized demonstrations yield significantly higher accuracy and power, with consistent improvement across optimization steps.
+3.3 Exploring Novel Optimization Strategies through Magnetic Field Adjustment
+When Few-shot ITR Performs Worse than Zero-shot ITR?
+LLMs exploit ITL by retrieving task demonstrations, yet the conditions under which demonstrations help remain unclear.
+We postulated that selecting the top-$k$ most semantically aligned examples would maximise model capacity.
+To test this idea we compared four retrieval policies: no demonstrations (Demo.), semantically irrelevant (no-meaning) demonstrations, bottom-$k$ and top-$k$ examples.
+Top-$k$ retrieval improved accuracy across tasks (Fig.
+6a).
+By contrast, no-meaning demonstrations, which exert a negative “semantic field”, reduced effective context utilisation and yielded performance below the zero-shot baseline.
+These observations show that the magnetic polarity of semantic alignment determines whether demonstrations aid ICL.
+The influence of diversity in ITL retrieval: a static vs. dynamic perspective.
+Two retrieval paradigms are widely employed: static retrieval, in which a fixed demonstration set is used for all test instances, and dynamic retrieval, in which demonstrations are chosen per instance.
+Semantic diversity enhances static retrieval by providing complementary contextual signals [23], whereas it can hinder dynamic retrieval by diluting alignment with the query [33].
+We quantified these effects with random retrieval, diversity-oriented static retrieval and similarity-oriented dynamic retrieval.
+Similarity-oriented dynamic retrieval achieved the highest accuracy (57.0%) and power-strength (>70) (Fig.
+6b).
+In static retrieval the same diverse set yielded higher accuracy than random retrieval, but in dynamic retrieval diversity conferred no benefit.
+Thus, diversity is advantageous only when demonstrations are reused across queries; when selection is instance-specific, alignment dominates.
+Demo-Optimization: a novel approach to enhance demonstration efficacy in ITL.
+Recently, gradient-based prompt optimisation can refine instructions for large language models [46].
+Building on this principle we introduce “demo-optimisation”, an algorithm that jointly adjusts the composition and ordering of demonstrations to balance their relative influence with the user prompt (See Methods in Appendix for more implementation details).
+As shown in Fig.
+6c, demo-optimisation increased accuracy by 20 percentage points relative to instruction-only optimisation and alleviated under-fitting in earlier approaches.
+The method therefore offers a scalable route to improve demonstration-driven reasoning.
+Figure 7: Optimization Exploration through Resistor Reorganization(a, b) and Unified Optimization for Applications (c, d) .
+a, Overview of the fine-grained self-consistency strategy in circuits.
+By decomposing parallel ITR resistors into smaller units, this approach effectively reduces overall circuit resistance.
+b, The Chain-of-Verification (CoV) strategy further optimizes circuit resistance by incorporating voting resistors into the self-consistency framework.
+This method reduces overall resistance by sampling and validating multiple voting results.
+c, Enhanced circuit performance is achieved by leveraging stronger magnetic fields and reducing logical resistance.
+The power-optimized model produced notable results in the IOI (left) and the IMO (right) tests.
+d, A methodology for integrating theoretical power into manual ratings during idea generation.
+This approach significantly improves the accuracy and reliability of the evaluation process compared to the original solution.
+3.4 Exploring Novel Optimization Strategies through Resistor Reorganization
+Optimising an equal proportion of model capabilities yields a performance gain that scales with the intrinsic difficulty of each capability.
+Reducing the resistance of the most challenging reasoning component from $R$ to $\frac{R}{k}$ therefore produces a larger decrease in the total inference-time resistance, $R_{\text{ITR}}$ , than an equivalent change applied to an easier component (See Theoretical Proof in Appendix for mathematical proof).
+Meta-analysis (Methods) of benchmark problems requiring long-horizon planning confirms this prediction: the exploration-centred reasoning procedure of model o1-preview [31] preferentially improves high-resistance components and delivers the largest gains on the most complex tasks.
+By contrast, tasks dominated by specialised domain knowledge benefit less, owing to limited optimisation of the corresponding knowledge component.
+Fine-grained self-consistency reduces verification resistance.
 Easily, we can prove that, with a stable supply voltage, we can split a self-consistency ensemble into more fine-grained step-by-step self-consistency by conducting majority voting verification for each step, so that each resistance value of verification resistor can be smaller, and we can obtain larger output power (see Theoretical Proof).
 Indeed, as the number of self-consistency steps increases, the theoretical power value gradually increases.
-As shown in Figure 6a, the corresponding performance also gradually increases.
+As shown in Fig.
+7a, the corresponding performance also gradually increases.
 At the same time, the corresponding upper bound and lower bound also satisfy the original self-consistency and coverage.
 This not only proves the effectiveness of our proposed fine-grained self-consistency method but also proves the practical usability of the theoretical model.
-Chain-of-Verification enhances model performance with parallelized verification resistors
-In conventional self-consistency frameworks, verification is typically conducted through series resistors.
-We propose an innovative approach, referred to as the “chain-of-verification,” which involves two key looped stages: first, verifying the correctness of individual outputs in parallel; and second, merging these verified results and subsequently meta-validating the correct to confirm the overall result of the merged outputs.
-It can enhance model performance by reducing the overall verification resistance.
-As depicted in Figure 6b, parallelization initially boosts performance by distributing verification tasks and lowering verification resistance.
-However, this improvement has limits.
-As parallelization increases, the overall resistors for parallel verifiers begins to rise.
-Beyond a critical threshold, this combined resistor surpasses the original self-consistency resistor, reversing the performance gains.
-As depicted in Figure 6b, this also demonstrates excessive chain-of-verification introduces diminishing returns.
-Thus, the chain-of-verification framework underscores the importance of balancing parallelization and resistor to optimize model verification efficiency.
-6 Best Practice
-To evaluate the impact of sufficient power on LLM performance, we analyzed its effects across three domains: the International Olympiad in Informatics (IOI), the International Mathematical Olympiad (IMO), and Natural Language Processing (NLP) idea generation using the AI-Scientist framework [23], by ensembling strategies outlined in Exploration Section.
-Our findings, shown in Figure 6c, demonstrate that augmenting the o1-preview model with enhanced power results in significant improvements.
-The power-enhanced o1-preview outperformed AlphaCode [20] and other state-of-the-art models on the IOI-level test set (code_contests [20]).
-Crucially, this boost consistently surpassed configurations lacking sufficient power, underscoring its pivotal role in model efficacy.
-In competitive contexts, such as IOI 2024 and IMO 2024, the powered o1-preview outperformed nearly 80% of global participants, earning silver and bronze medals.
-This highlights power as a decisive factor in high-stakes scenarios where traditional optimizations often fall short.
-Beyond correctness-focused tasks, as shown in Figure 6d, sufficient power also enhanced performance in creative challenges.
-For instance, in NLP-related idea-generation, it facilitates the better generation of novel and impactful research ideas (over 1 point of overall score improvement on average), expanding the potential of LLMs in both objective and subjective tasks.
+Chain-of-verification balances parallel and serial checks.
+Traditional self-consistency frameworks rely on a sequential verification, where each step depends on the completion of the previous one.
+This structure often incurs high computational costs, as subsequent tasks cannot begin until earlier ones are resolved.
+To overcome this limitation, we introduce the “chain-of-verification (CoV)," a novel strategy that parallelizes verification tasks into independent components.
+Each component operates concurrently, validating distinct parts of the system without interdependence.
+This parallel approach reduces overall computational resistance by distributing the workload across multiple verifiers, thereby enabling more effective validation.
+As illustrated in Fig.
+7b, parallelization not only accelerates the verification process but also enhances model performance by mitigating bottlenecks.
+This improvement is particularly pronounced in large-scale systems, where sequential verification would otherwise impede efficiency.
+The effectiveness of CoV can also be mathematically proved in Theoretical Proof.
+3.5 Unified Optimization for Downstream Applications
+Adequate computational power is a primary determinant of LLM performance across disparate tasks.
+We benchmarked an enhanced version of the o1-preview model on algorithmic problem solving (International Olympiad in Informatics, IOI), mathematical reasoning (International Mathematical Olympiad, IMO) and open-ended idea generation in the AI-Scientist framework [28], using the ensemble procedure described above.
+Increasing the output power yielded marked gains (Figure 7c): on the IOI code_contests [24] test set the power-augmented model outperformed AlphaCode [24] and all other published baselines.
+In live competitions (IOI 2024 and IMO 2024) it surpassed 8̃0% of human participants, earning silver and bronze medals, respectively, and highlighting compute as a critical factor when conventional optimisation is insufficient.
+Enhanced compute also improved performance in creativity-oriented tasks.
+In the AI-Scientist benchmark, the same configuration increased the mean expert-rated novelty and impact scores by >1.0 (Fig.
+7d), indicating that scaling compute extends the applicability of LLMs beyond correctness-centred objectives.
 4 Conclusion & Discussion
-The Electronic Circuit Model (ECM) offers a unified framework for understanding and optimizing the ICL and CoT mechanisms, marking a significant advancement in LLM research.
-This framework elucidates model behavior while providing a quantifiable approach to enhance LLM performance.
-Our experiments demonstrate that ECM effectively captures model intricacies, serving as a reliable tool for analyzing and improving AI capabilities.
-Moreover, applying power-based optimization derived from ECM to advanced reasoning challenges, such as the International Olympiad in Informatics and the International Mathematical Olympiad, yielded superior results compared to leading competitors.
-These outcomes validate the theoretical foundation of ECM and highlight its transformative potential in improving AI reasoning and learning efficiency, with broad implications for next-generation intelligent systems.
+The Electronic Circuit Principles (ECP) introduces a unified principle for understanding and optimizing LLMs’ performance.
+By quantifying internal dynamics, ECP clarifies model behavior and guides modular performance improvements.
+Our experiments confirm that ECP reliably captures model complexities, providing a robust tool for AI analysis and enhancement.
+Moreover, ECP-driven power optimizations applied to the International Olympiad in Informatics and the International Mathematical Olympiad outperform current leading methods.
+These findings validate ECP’s theoretical foundation and underscore its potential to elevate AI reasoning and learning efficiency in next-generation systems.
 5 Code Availability
-The code and relevant data are available at https://github.com/LightChen233/ECM.
-A Method Implementation Details
-A.1 Theoretical Model
-Faraday’s Law of In-Context Learning
-In-context learning (ICL) can be likened to the behavior of a semantic magnetic field generated by a sub-power source, as described by Faraday’s Law.
+The code and relevant data are available at https://github.com/LightChen233/ECP.
+A Methods
+A.1 Electronic Circuit Principles (ECP)
+Faraday’s Law of Inference-Time Learning
+Inference-time learning (ITL), also named as in-context learning, can be likened to the behavior of a semantic magnetic field generated by a sub-power source, as described by Faraday’s Law.
 The positive direction of the semantic magnetic field is defined by the query semantic vector $\overrightarrow{S}_{q}$.
@@
 where $\lambda$ is the decay constant representing the rate at which the model forgets or absorbs information.
-According to Faraday’s Law of electromagnetic induction, the rate of change in the semantic magnetic field induces an electromotive force, which drives effective ICL.
-With a time step $\Delta t$ for each reasoning step, the induced EMF, $\mathcal{E}_{\text{ICL}}$, is:
-$\mathcal{E}_{\text{ICL}}=-\frac{\Delta\Phi^{B}(t)}{\Delta t}=\lambda\Phi^{B}_{0}.$ (6)
+According to Faraday’s Law of electromagnetic induction, the rate of change in the semantic magnetic field induces an electromotive force, which drives effective ITL.
+With a time step $\Delta t$ for each reasoning step, the induced EMF, $\mathcal{E}_{\text{ITL}}$, is:
+$\mathcal{E}_{\text{ITL}}=-\frac{\Delta\Phi^{B}(t)}{\Delta t}=\lambda\Phi^{B}_{0}.$ (7)
 Substituting $\Phi^{B}_{0}$ into this equation yields:
-$\mathcal{E}_{\text{ICL}}=\lambda\sum^{N}_{i=1}\frac{\overrightarrow{S}_{q}\cdot\overrightarrow{S}_{i}}{\left\lvert\overrightarrow{S}_{q}\right\rvert}.$ (7)
-In conclusion, this analogy offers a structured framework to analyze how information flows during ICL.
-By extending the dynamics of ICL through the principles of electromagnetic induction, we provide a mathematical model to explain how models incorporate and gradually forget information as they process contextual examples.
-Ohm’s Law of Chain-of-Thought
+$\mathcal{E}_{\text{ITL}}=\lambda\sum^{N}_{i=1}\frac{\overrightarrow{S}_{q}\cdot\overrightarrow{S}_{i}}{\left\lvert\overrightarrow{S}_{q}\right\rvert}$. (8)
+In conclusion, this analogy offers a structured framework to analyze how information flows during ITL.
+By extending the dynamics of ITL through the principles of electromagnetic induction, we provide a mathematical principle to explain how models incorporate and gradually forget information as they process contextual examples.
+Ohm’s Law of Inference-Time Reasoning
 In this framework, each sub-difficulty of a model introduces a distinct level of difficulty, analogous to resistors in series.
@@
 [4], the ease with which a model can solve a particular task is represented by its reasoning boundary.
-Let $\mathcal{B}_{\text{CoT}}$ denote the overall reasoning boundary of a Chain-of-Thought process, and $\mathcal{B}_{i}$ denote the reasoning boundary of the $i$-th sub-capability.
+Let $\mathcal{B}_{\text{ITR}}$ denote the overall reasoning boundary of a Inference-Time Reasoning process, and $\mathcal{B}_{i}$ denote the reasoning boundary of the $i$-th sub-capability.
 The overall reasoning boundary can be expressed by the harmonic sum:
-$\mathcal{B}_{\text{CoT}}=\frac{1}{\sum^{K}_{i=1}\frac{1}{\mathcal{B}_{i}}},$ (8)
+$\mathcal{B}_{\text{ITR}}=\frac{1}{\sum^{K}_{i=1}\frac{1}{\mathcal{B}_{i}}},$ (9)
 where $K$ denotes the number of sub-capabilities.
 Reasoning resistor, $R$, quantifies the difficulty of a reasoning process and is mathematically defined as the reciprocal of the reasoning boundary.
-A higher boundary corresponds to an easier problem and lower resistor.
+A higher boundary corresponds to an easier problem and lower resistance.
 Specifically, we define the sub-difficulty of the $i$-th sub-task as:
 $R_{i}=\frac{1}{\mathcal{B}_{i}}.$ (10)
-Substituting this into Equation 8, the total reasoning resistor for the CoT process is expressed as:
-$R_{\text{CoT}}=\sum^{K}_{i=1}R_{i},$ (10)
-where $R_{\text{CoT}}$ represents the combined difficulty, analogous to total resistor in a series circuit. Overall, the CoT reasoning framework is further formalized by the Ohm’s law analogy:
-$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}}{R_{\text{CoT}}+R_{0}},$ (11)
-where $R_{0}$ denotes a output resistor with static resistance value. This modular formulation effectively quantifies cognitive task complexity through combined reasoning resistor.
-A.2 Genral Experimental Settings
+Substituting this into Equation 9, the total reasoning resistor for the ITR process is expressed as:
+$R_{\text{ITR}}=\sum^{K}_{i=1}R_{i},$ (11)
+where $R_{\text{ITR}}$ represents the combined difficulty, analogous to a total resistor in a series circuit. Overall, the ITR reasoning framework is further formalized by the Ohm’s law analogy:
+$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}}{R_{\text{ITR}}+R_{0}},$ (12)
+where $R_{0}$ denotes an output resistor with static resistance value. This modular formulation effectively quantifies cognitive task complexity through a combined reasoning resistor.
+A.2 General Experimental Settings
 All experiments, unless stated otherwise, are conducted on GPT-3.5-Turbo.
 Following Wei et al.
-[41], CoT experiments select 5 manually constructed demonstrations for all multi-step reasoning tasks.
+[43], ITR experiments select 5 manually constructed demonstrations for all multi-step reasoning tasks.
 Top-p is selected from $\{0.95,1\}$, and Temperature is selected in range $[0,1]$ for robustness evaluation.
-Experiments are primarily conducted on the BigGSM dataset [5], a benchmark for mathematical reasoning.
-In ICL, demonstrations are dynamically retrieved using a 5-shot approach from the GSM8K development dataset [8] by default.
 To formalize the demonstration selection process, we retrieve the semantic magnetic field strength $\Phi^{B}_{0}$ by default, as defined in Eq.
-4.
-All demonstration representation methods employ the meta-recognition [10] framework by default: GPT-4o [1] extracts sub-capability tags, and Roberta [22] concatenates and represents these tags in the hidden space.
-In addition, only the magnetic field strength $\Phi^{B}_{0}$ is calculated and the resistor value $R_{\text{CoT}}$ directly adopts the fitting values from Chen et al.
+5.
+In addition, only the magnetic field strength $\Phi^{B}_{0}$ is calculated and the resistor value $R_{\text{ITR}}$ directly adopts the fitting values from Chen et al.
 [4].
-In addition, other parameters are fitted values using the verification set.
+Further, other parameters are fitted values using the verification set.
 Different models fit different voltages $\mathcal{E}_{model}$.
@@
 All experiments involving variance were sampled 3 times at different temperatures.
-A.3 Model Verification Experiment Details
-CoT Ohm’s Law Verification
-To validate the Ohm’s Law assumption in the CoT approach, we conduct the following experiment.
-To eliminate potential interference from ICL and ensure effective stimulation of CoT capabilities, we append the phrase “Let’s think step-by-step!" to the user query.
-This adjustment aims to stimulate and isolate the CoT mechanism and prevent ICL-related influences.
-We then perform sampling based on power levels from the BigGSM dataset, using 30-unit intervals to map calculated power to accuracy [^1] The absolute power values are meaningless due to inconsistencies in semantic space scaling across encoding methods.
-However, their relative values within a method correlate positively with real accuracy..
+Notably, in addition to the validation experiments, we predominantly utilize the BigGSM [4] and BigGSM++ [5] datasets, which precisely annotate reasoning difficulty across global planning, local operations, and calculation sub-tasks.
+In ITL, demonstrations are dynamically retrieved using a 5-shot approach from the GSM8K development dataset [8] by default.
+A.3 Verification Experiment Details
+In the effectiveness validation section of ECP, and with the aim of comprehensively assessing the efficacy of our tasks, we first classify each task’s difficulty (hereafter “resistance”) into four categories: global planning difficulty, local operational difficulty, domain‐knowledge analysis difficulty, and computational difficulty (the latter being relevant only to mathematical tasks).
+The total resistance is then defined as:
+$R_{\text{ITR}}=R_{\text{plan}}+R_{\text{operation}}+R_{\text{domain}}+R_{\text{calculate}},$ (13)
+where $R_{\text{plan}}$ denotes the resistance due to global planning difficulty, $R_{\text{operation}}$ denotes the resistance due to local operational difficulty, $R_{\text{domain}}$ denotes the resistance due to domain‐knowledge analysis difficulty, and $R_{\text{calculate}}$ denotes the resistance due to computational difficulty (set to zero for non‐computational tasks).
+Furthermore, as illustrated in the Fig.
+8, the task evaluation benchmarks are primarily drawn from the BigGSM [4], BigGSM++ [5], MedProbing [7], Med-MCQA [32], MedQA [18], PubMedQA [19], BBH [38], MMLU [16], MMLU-Pro [42], GPQA [36], and SuperGPQA [12] datasets with over 350 tasks.
+[…]
+The annotation of the number of global and local reasoning steps required depends on the model-generated correct rationales and is automated through a manually constructed step-segmentation annotation program.
+Furthermore, since the granularity of the model’s reasoning does not fully align with that of humans, and the segmentation program cannot perform fine-grained annotations for every case, a considerable amount of annotation noise exists in the validation experiments across nearly 350 tasks, resulting in a relatively low $R^{2}$.
+However, under a strictly human-annotated setting for global and local step counts, such as that of the BigGSM++ dataset, the $R^{2}$ can exceed $0.9$.
+Therefore, except for the validation experiments, all subsequent experiments are conducted on the BigGSM++ dataset.
+Figure 8: Over 350 tasks in unified evaluation data.
+Inspired by Chen et al.
+[4], we treat the domain difficulty as a constant for each task; this constant was obtained via parameter search using Newton’s method in a validation experiment on 10% of the GPT-3.5-turbo generated dataset.
+For mathematical tasks, we design task-specific regular-expression extraction rules to estimate computational difficulty; however, since most operations not involving large numbers yield a computational difficulty close to zero, we omit this factor for some math tasks.
+ITR Ohm’s Law Verification
+To validate the Ohm’s Law assumption in the ITR approach, we conduct the following experiment.
+To eliminate potential interference from ITL and ensure effective stimulation of ITR capabilities, we append the phrase “Let’s think step-by-step!" to the user query.
+This adjustment aims to stimulate and isolate the ITR mechanism and prevent ITL-related influences.
+We then perform sampling based on power levels from the BigGSM dataset, using task-specific unit intervals to map calculated power to accuracy [^1] Since the distribution of reasoning difficulty varies across tasks, a sufficiently large sample size is required to accurately compute accuracy.
+Accordingly, we employ different units for each task..
 To reduce random errors and ensure sufficient sampling points, we retain data points with at least 10 samples.
-Accuracy is calculated for each sampling interval to analyze the relationship between power levels and model performance, enabling an effective evaluation of the correlation between sampling power and accuracy.
-ICL Faraday’s Law Verification
-To validate the semantic magnetic field hypothesis in ICL, we observe that without CoT reasoning, performance on complex mathematical tasks often declines to near-zero levels, thereby cannot easily observe their relevance.
-Thus, this experiment incorporates both CoT reasoning and ICL.
-While the CoT reasoning format remains constant, we focus on varying the similarity measures used for sample retrieval through ICL.
+Accuracy is calculated for each sampling interval to analyze the relationship between power levels and model performance, enabling an effective evaluation of the correlation between output power and accuracy to verify Ohm’s Law.
+Figure 9: The Analysis for ITL Faraday’s Law Verification Across 350 Tasks and 4 Models.
+ITL Faraday’s Law Verification
+To validate the semantic magnetic field hypothesis in ITL, we observe that without ITR reasoning, performance on complex mathematical tasks often declines to near-zero levels, thereby cannot easily observe their relevance.
+Thus, this experiment incorporates both ITR reasoning and ITL.
+While the ITR reasoning format remains constant, we focus on varying the similarity measures used for sample retrieval through ITL.
 The central adjustment involves applying different similarity measures for sample retrieval, directly affecting the model’s performance on complex tasks.
@@
 All demonstration representation methods employ BGE [3] to represent the request in the hidden space.
-Unified circuit mechanism verification
-To further validate the effectiveness of our unified framework for modeling ICL and CoT, we conduct experiments integrating both paradigms into different retrieval settings.
-These experiments test the robustness and rationality of the proposed theoretical framework.
-Specifically, we examine three retrieval methods: random retrieval, forward magnetic field retrieval, and reverse magnetic field retrieval.
-In the random retrieval setup, samples are selected randomly from the development set as ICL demonstrations for each test sample.
-The Retrieval-ICL selects samples with maximized magnetic field intensity in the corresponding test sample’s direction.
-Conversely, the Reverse-ICL focuses on samples with minimal (even negative) magnetic field intensity, enabling exploration of contrasting ICL examples.
-This comprehensive validation highlights the robustness of our framework across diverse retrieval strategies and empirically demonstrates its generalizability in both ICL and CoT scenarios.
-Extension Experiment Details on Various Models and Tasks
-To further enhance the applicability of ECM, we conducted experiments across a series of models, where each model was designed to fit specific baseline voltage values, $\Phi^{B}_{0}$, and compute the corresponding power output.
-Subsequently, all models were consolidated onto a single graph, enabling us to analyze the correlation between output power and relative accuracy.
-As illustrated in Figure x, the results reveal clear and detailed correlations across the models.
-However, when accuracy exceeds 90%, the scarcity of samples requiring sufficient power in this range, likely due to limitations in the dataset, leads to poor fitting in this region.
-Furthermore, our experiments on HotpotQA and Medical Probing closely adhered to the experimental settings established by Chen et al.
-[5].
-However, since Medical Probing includes rationale annotations limited to background knowledge rather than chain reasoning, we supplemented portions of the reasoning process to facilitate a more precise correlation analysis.
-Additionally, the data for robotic planning was directly extracted from the running logs provided by Valmeekam et al.
-[34].
-A.4 Model Explanation Experiments
-A.4.1 ICL Explanation
-Embedder Comparison
-To evaluate how different embedders influence model performance, we replace the sample encoder with BERT [9], RoBERTa [22], and BGE [4].
-These embeddings are used for sample retrieval in ICL to minimize potential biases that may arise if a single embedder disproportionately emphasizes specific data features.
-For instance, an embedder might favor samples aligned with its encoding patterns, leading to distortions in the relationship between retrieved representations and model performance.
-By employing diverse embedders, we aim to mitigate such biases and promote more accurate theoretical modeling for ICL.
+A.4 Optimization Interpreting Experiments through Magnetic Field Adjustment
+Encoder Comparison
+To evaluate how different encoders influence model performance, we replace the sample encoder with BERT [9], RoBERTa [27], and BGE [3].
+These embeddings are used for sample retrieval in ITL to minimize potential biases that may arise if a single encoder disproportionately emphasizes specific data features.
+For instance, an encoder might favor samples aligned with its encoding patterns, leading to distortions in the relationship between retrieved representations and model performance.
+By employing diverse encoders, we aim to mitigate such biases and promote more accurate theoretical modeling for ITL.
+Meta-Recognition Representations
+In this experiments, all demonstration representation methods employ the meta-recognition [10] framework by default.
+Specifically, GPT-4o [1] is utilized to extract sub-capability tags, and then RoBERTa [27] concatenates and represents these tags in the hidden space.
 Synthetic-CoT Implementation
-This section examines the implementation of Synthetic-CoT within the ECM framework to assess its impact on reasoning performance.
-Synthetic-CoT autonomously generates in-context demonstrations with chain-of-thought reasoning paths.
-We hypothesize that these demonstrations align the system’s magnetic field, reducing logical resistor and enhancing overall circuit power output.
-To validate this, we generate a series of demonstrations with two categories by Synthetic-CoT, including:
-  • Positive Magnetic Field Demonstrations: In-context demonstrations generated by Synthetic-CoT are positively aligned with the magnetic polarity of user queries direction.
-  • Negative Magnetic Field Demonstrations: In-context demonstrations generated by Synthetic-CoT are with opposing magnetic polarity of user queries direction.
+This section examines the implementation of Synthetic-CoT within the ECP framework to assess its impact on reasoning performance.
+Synthetic-CoT autonomously generates in-context demonstrations with Inference-Time Reasoning reasoning paths.
+We hypothesize that these demonstrations align the system’s magnetic field, reducing logical resistance and enhancing overall circuit power output.
+To validate this, we generate a series of demonstrations by Synthetic-CoT.
+Specifically, In-context demonstrations generated by Synthetic-CoT are positively aligned with the magnetic polarity of user queries direction.
 We analyze the magnetic field contributions of these categories, focusing on their differing effects.
 Subsequently, we integrate the computed power and measured accuracy to further investigate their operational mechanisms.
-A.4.2 CoT Explanation
+A.5 Optimization Interpreting Experiments through Resistor Reorganization
 Direct answer output
-We analyze the effect of direct answer output on reasoning performance through the ECM framework.
-This approach bypasses intermediary reasoning steps by appending the directive, “Let’s directly output the result without additional reasoning content,” which generates significant logical resistor within the system.
-To quantify this, we gradually improve the resistance value of reasoning resistor $R_{\text{CoT}}$, which meets largest Spearman correlations in the validation set, and employ the $R_{\text{CoT}}$ to the test set.
+We analyze the effect of direct answer output on reasoning performance through the ECP framework.
+This approach bypasses intermediary reasoning steps by appending the directive, “Let’s directly output the result without additional reasoning content,” which generates a significant logical resistor within the system.
+To quantify this, we gradually improve the resistance value of reasoning resistor $R_{\text{ITR}}$, which meets the largest Spearman correlations in the validation set, and employ the $R_{\text{ITR}}$ to the test set.
 Tool-Usage and PoT
-We further investigate how Tool-Usage and Program-of-Thought (PoT) mitigate specific resistors in the reasoning process, focusing on their impact under the ECM framework on computational and planning resistors:
+We further investigate how Tool-Usage and Program-of-Thought (PoT) mitigate specific resistors in the reasoning process, focusing on their impact under the ECP framework on computational and planning resistors:
   • Tool-Usage: Tool-Usage aims to reduce computational resistor by offloading complex calculations to external tools.
@@
 [^2] All implementation prompts come from Chen et al.
-[5]
-These findings confirm that while Tool-Usage primarily addresses single-step computational resistor, PoT excels in optimizing overall planning resistor.
-Both methods significantly enhance performance within the ECM framework, highlighting their complementary contributions to reducing resistor and improving reasoning capabilities.
+[4].
+These findings confirm that while Tool-Usage primarily addresses the single-step computational resistor, PoT further excels in optimizing the overall planning resistor.
+Both methods significantly enhance performance within the ECP framework, highlighting their complementary contributions to reducing resistance and improving reasoning capabilities.
 Self-consistency and Coverage Design
@@
 [41], which employs multiple parallel reasoning paths and a majority-vote mechanism for the final decision.
-This strategy is hypothesized to enhance performance by reducing the system’s resistance to diverse inputs, analogous to parallel resistors in a circuit, where total resistor decreases compared to series connections.
+This strategy is hypothesized to enhance performance by reducing the system’s resistance to diverse inputs, analogous to parallel resistors in a circuit, where the total resistance decreases compared to series connections.
 By applying this analogy, we demonstrate how parallel reasoning can lower cognitive load, improving accuracy and efficiency in decision-making.
 To examine the influence of inference scaling on model performance, we investigated the relationship between the number of parallel samples, $k$, and the metric like Accuracy or Pass@k, as described in the Inference Scaling Law.
-Specifically, Pass@k measures the accuracy of Coverage method which obtains a correct result when at least one sampled prediction is accurate.
+Specifically, Pass@k measures the accuracy of the Coverage method which obtains a correct result when at least one sampled prediction is accurate.
 We modeled the reasoning process as a resistive circuit, where each sample represents a parallel resistor.
@@
 Further, Coverage methods introduce zero verification resistance introduce much less overall resistance values, larger output power, and higher related performance.
-The experiment systematically varied $k$ from 1 to 100 in increments of 5, recording the accuracy and Pass@k metric under each condition [^3] We remove those generated samples are totally the same to keep the independent sampling..
+The experiment systematically varied $k$ from 1 to 100 in increments of 5, recording the accuracy and Pass@k metric under each condition.
 We track individual reasoning path resistance, comparing these values with observed accuracy to validate the circuit analogy.
-The effectiveness of Self-consistency and Coverage methods can be mathematically proved based on the ECM detailed in Theoretical Proof.
-A.5 Model Exploration Experiments
-A.5.1 ICL Exploration
+The effectiveness of Self-consistency and Coverage methods can be mathematically proved based on the ECP detailed in Theoretical Proof.
+We annotated 100 instances, each comprising 32 outputs sampled from the model, with four annotators involved in the labeling.
+For each instance, the 32 outputs were presented to the annotators in sequence; they were asked to indicate whether each output contained text or logical content that was exactly duplicated from any previously shown example.
+Finally, for sample sizes of $2^{n}$ , we counted the number of outputs exhibiting no duplication and treated these counts as independent samples.
+A.6 Optimization Exploration Experiments through Magnetic Field Adjustment
 Diversity
-This study examines the differing impacts of semantic diversity on static and dynamic ICL retrieval.
+This study examines the differing impacts of semantic diversity on static and dynamic ITL retrieval.
 We map all samples into the hidden space using RoBERTa.
-In static ICL, we employ a fixed set of demonstrations with the highest semantic diversity across the task, hypothesizing that this diversity enhances model performance by providing a wider range of contextual cues.
-In contrast, dynamic ICL selects demonstrations in real-time based on the test sample’s characteristics, prioritizing semantically similar examples that align closely with the input.
+In static ITL, we employ a fixed set of demonstrations with the highest semantic diversity across the task, hypothesizing that this diversity enhances model performance by providing a wider range of contextual cues.
+In contrast, dynamic ITL selects demonstrations in real-time based on the test sample’s characteristics, prioritizing semantically similar examples that align closely with the input.
 To test these, we evaluate both retrieval strategies.
-For static ICL, we compare models exposed to high-diversity and random demonstration sets, measuring performance using accuracy.
-For dynamic ICL, we contrast random selection with retrieval strategies emphasizing semantic alignment.
-We hypothesize that semantic diversity improves static ICL by stabilizing decision-making, while in dynamic ICL, retrieving highly relevant demonstrations yields better accuracy.
-In addition, in order to prove that diversity is ineffective in the setting of dynamic ICL, we sort the semantic magnetic field strength and select as diverse samples as possible among the top-$k$ samples as ICL demonstrations.
+For static ITL, we compare ECP exposed to high-diversity and random demonstration sets, measuring performance using accuracy.
+For dynamic ITL, we contrast random selection with retrieval strategies emphasizing semantic alignment.
+We hypothesize that semantic diversity improves static ITL by stabilizing decision-making, while in dynamic ITL, retrieving highly relevant demonstrations yields better accuracy.
+In addition, in order to prove that diversity is ineffective in the setting of dynamic ITL, we sort the semantic magnetic field strength and select as diverse samples as possible among the top-$k$ samples as ITL demonstrations.
 Demo-Optimization
-This experiment introduces demo-optimization, a novel method for optimizing demonstrations in ICL.
+This experiment introduces demo-optimization, a novel method for optimizing demonstrations in ITL.
 The approach dynamically refines demonstrations based on feedback from the LLM’s outputs, which serve as gradients for improvement.
-Specifically, initial ICL demonstrations in prompting are selected, with both development dataset for optimization and optimization prompt configurations prepared.
+Specifically, initial ITL demonstrations in prompting are selected, with both development dataset for optimization and optimization prompt configurations prepared.
 A feedback loop guides adjustments to the demonstrations, improving their clarity, relevance, and alignment with prompt requirements.
@@
 The method is compared against static and whole-prompt-optimized baselines, with performance assessed using task accuracy, output consistency, and learning efficiency.
-A.5.2 CoT Exploration
-CoT Optimization Proportion
+A.7 Optimization Exploration Experiments through Resistor Reorganization
+ITR Optimization Proportion
 Figure 10: Meta-analysis for the model improvements on different tasks.
@@
 These datasets are divided into two main categories: (1) tasks requiring advanced planning and logical reasoning, such as mathematical derivations and multi-hop question answering, and (2) tasks reliant on extensive domain-specific knowledge, such as chemistry and physical question answering.
-As shown in Figure 7, our meta-analysis demonstrates that tasks requiring advanced planning gain the most from enhanced exploratory reasoning, significantly improving challenging components (e.g., a 2.8-fold optimization in planning capability) and overall task performance (e.g., a 3.6-fold optimization on AIME, a benchmark emphasizing long-term planning).
+As shown in Fig.
+10, our meta-analysis demonstrates that tasks requiring advanced planning gain the most from enhanced exploratory reasoning, significantly improving challenging components (e.g., a 2.8-fold optimization in planning capability) and overall task performance (e.g., a 3.6-fold optimization on AIME, a benchmark emphasizing long-term planning).
 In contrast, tasks relying heavily on domain-specific knowledge exhibit limited improvement without targeted optimization in those areas.
@@
 For example, tasks involving complex reasoning and knowledge integration, such as GPQA (chemistry, physics) and HotpotQA, show performance improvements directly proportional to advancements in the related knowledge domains (approximately 1:1 ratio).
-The effectiveness and the underlying reasons for such phenomena can be mathematically analyzed and proved based on the ECM detailed in Theoretical Proof.
+The effectiveness and the underlying reasons for such phenomena can be mathematically analyzed and proved based on the ECP detailed in Theoretical Proof.
 Fine-grained Self-consistency
-In this experiment, we examine how decomposing a self-consistency ensemble into smaller, incremental majority-voting steps reduces reasoning resistor in circuits, compared to the conventional single-step approach.
+In this experiment, we examine how decomposing a self-consistency ensemble into smaller, incremental majority-voting steps reduces the reasoning resistance in circuits, compared to the conventional single-step approach.
 The circuit operates at a stable supply voltage of 5 units, and reasoning resistance value and output power are evaluated as the number of steps increases.
@@
 Our findings demonstrate a consistent decline in reasoning resistance alongside a rise in output power, confirming that finer-grained majority-voting processes improve both efficiency and power output in circuits.
-The effectiveness of this method can be mathematically proved based on the ECM detailed in Theoretical Proof.
+The effectiveness of this method can be mathematically proved based on the ECP in Theoretical Proof.
 Chain-of-Verification
-Traditional self-consistency frameworks rely on a sequential verification model, where each step depends on the completion of the previous one.
+Traditional self-consistency frameworks rely on a sequential verification, where each step depends on the completion of the previous one.
 This structure often incurs high computational costs, as subsequent tasks cannot begin until earlier ones are resolved.
@@
 This parallel approach reduces overall computational resistance by distributing the workload across multiple verifiers, thereby enabling more effective validation.
-As illustrated in Figure 6b, parallelization not only accelerates the verification process but also enhances model performance by mitigating bottlenecks.
+As illustrated in Fig.
+7b, parallelization not only accelerates the verification process but also enhances model performance by mitigating bottlenecks.
 This improvement is particularly pronounced in large-scale systems, where sequential verification would otherwise impede efficiency.
-The effectiveness of CoV can also be mathematically proved based on the ECM detailed in Theoretical Proof.
-A.6 Best Practice
-In practice, we integrate all proposed strategies into the O1-preview model to address IOI, IMO problem-solving, and idea-generation tasks for much less resistance value of planning resistor.
+The effectiveness of CoV can also be mathematically proved in Theoretical Proof.
+A.8 Best Practice
+In practice, we integrate all proposed strategies into the o1-preview model to address IOI, IMO, and idea-generation tasks for much less resistance value of planning resistor.
 IOI Task
-For the IOI task, we employ the following methods: First, using O1-preview, we maximize the optimization of the resistance value of its “planning resistor." Additionally, five high-intensity semantic field problems from the code_contest training dataset are retrieved as demonstrations, boosting the voltage supplied by additional ICL fields.
-Second, we apply the Coverage technique, sampling 100 instances and leveraging both provided and model-generated test cases in parallel, effectively reducing resistance in chain-of-thought reasoning.
-This process, executed in parallel, effectively reduces the resistance in the chain-of-thought reasoning.
+For the IOI task, we employ the following methods: First, using o1-preview, we maximize the optimization of the resistance value of its “planning resistor." Additionally, five high-intensity semantic field problems from the code_contest training dataset are retrieved as demonstrations, boosting the voltage supplied by additional ITL fields.
+Second, we apply the Coverage technique, sampling 100 instances and leveraging both provided and model-generated test cases in parallel, effectively reducing resistance in Inference-Time Reasoning reasoning.
+This process, executed in parallel, effectively reduces the resistance in the Inference-Time Reasoning reasoning.
 Furthermore, we incorporate the CoV concept to iteratively validate the final outputs, thereby optimizing the final performance.
@@
 A similar approach is adopted for the IMO task.
-O1-preview is optimized for planning resistor, and five high-intensity semantic field problems from past IMO datasets are retrieved as demonstrations to enhance ICL field voltage.
+o1-preview is optimized for planning resistor, and five high-intensity semantic field problems from past IMO datasets are retrieved as demonstrations to enhance ITL field voltage.
 The Coverage technique is again employed, sampling 100 instances in parallel to reduce reasoning resistance.
@@
 Each paper is reviewed by at least four reviewers, including two Senior or Superior reviewers, following NeurIPS guidelines.
-High-quality examples with strong semantic fields, sourced from past ICLR, NeurIPS, and CoLM openreview data, are used to guide generation.
+High-quality examples with strong semantic fields, sourced from past ITLR, NeurIPS, and CoLM openreview data, are used to guide generation.
 The Coverage technique, sampling of 100 instances, and iterative validation through the CoV framework further refine the process.
-Figure 8: Neuron activation heatmaps for the most activated attention layer under three reasoning strategies: (a) Zero-shot CoT, (b) ICL with Semantic Magnetic Field Retrieval, and (c) Program of Thought (PoT).
-A.7 Internal Mechanism underlying the ECM
-To investigate the internal mechanisms underlying ECM, we conduct experiments in the LLaMA3-8B [12] focusing on three strategies: Zero-shot CoT, ICL with Semantic Magnetic Field Retrieval, and PoT.
-By analyzing neuron activation values and frequencies—particularly in the most activated attention layer—we aim to elucidate the relationship between reasoning strategies and neural dynamics.
-These findings support the theoretical framework of ECM.
+Figure 11: Neuron activation heatmaps for the most activated attention layer under three reasoning strategies: (a) Zero-shot ITR, (b) ITL with Semantic Magnetic Field Retrieval, and (c) Program of Thought (PoT).
+A.9 Internal Mechanism underlying the ECP
+To investigate the internal mechanisms underlying ECP, we conduct experiments in the LLaMA3-8B [13] focusing on three strategies: Zero-shot ITR, ITL with Semantic Magnetic Field Retrieval, and PoT.
+By analyzing neuron activation values and frequencies, particularly in the most activated attention layer, we aim to elucidate the relationship between reasoning strategies and neural dynamics.
+These findings support the theoretical framework of ECP.
 The experiments use the BigGSM dataset, a standard benchmark for mathematical reasoning tasks.
-For ICL, demonstrations are dynamically retrieved using the BGE model in a 3-shot configuration.
-Figure 8 shows neuron activation heatmaps for the most activated attention layer under each reasoning strategy.
-For Zero-shot CoT, the heatmap reveals sparse activations, with limited regions of intense activity, highlighting the difficulty of reasoning without contextual support.
-Conversely, ICL and PoT display more distributed and elevated activation patterns.
+For ITL, demonstrations are dynamically retrieved using the BGE model in a 3-shot configuration.
+Fig.
+11 shows neuron activation heatmaps for the most activated attention layer under each reasoning strategy.
+For Zero-shot ITR, the heatmap reveals sparse activations, with limited regions of intense activity, highlighting the difficulty of reasoning without contextual support.
+Conversely, ITL and PoT display more distributed and elevated activation patterns.
 These suggest that the underlying reasoning mechanisms engage a broader spectrum of neurons, with comparable activation dynamics reflecting similar reasoning power across these strategies.
@@
 B.1.1 Analysis of Effectiveness of Self-consistency
-First, according to the series resistor assumption of CoT, the total resistance of series resistors can be expressed as:
-$R_{\text{all}}=R_{0}+R_{\text{CoT}}.$ (12)
-Building upon the illustration in Figure 4c, we hypothesize that self-consistency, achieved through multiple parallel reasoning processes and subsequent result aggregation by majority voting, can be represented as an aggregation resistor integrated alongside several parallel CoT resistors.
-In this framework, the effective resistance of the reasoning process transitions from $R_{\text{CoT}}$ to $\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}$, accompanied with a majority-voting resistor $R_{S}$ to determine the final outcome.
+First, according to the series resistor assumption of ITR, the total resistance of series resistors can be expressed as:
+$R_{\text{all}}=R_{0}+R_{\text{ITR}}.$ (14)
+Building upon the illustration in Fig.
+5c, we hypothesize that self-consistency, achieved through multiple parallel reasoning processes and subsequent result aggregation by majority voting, can be represented as an aggregation resistor integrated alongside several parallel ITR resistors.
+In this framework, the effective resistance of the reasoning process transitions from $R_{\text{ITR}}$ to $\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}$, accompanied with a majority-voting resistor $R_{S}$ to determine the final outcome.
 Consequently, the total resistance associated with self-consistency, denoted as $R^{{}^{\prime}}_{\text{all}}$, is expressed as:
-$R^{{}^{\prime}}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}},$ (13)
-where $n$ is the number of samples, $R_{\text{CoT}}^{i}$ denotes the resistance corresponding to the $i-$th sample and $R_{S}$ accounts for the resistor due to self-consistency integration.
+$R^{{}^{\prime}}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}},$ (15)
+where $n$ is the number of samples, $R_{\text{ITR}}^{i}$ denotes the resistance corresponding to the $i$-th sample and $R_{S}$ accounts for the resistor due to self-consistency integration.
 This framework allows for rapid computation of the output power:
-$P^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}})^{2}}.$ (14)
-As the number of samples increases, the curves in Figure 4c reveal that the real performance curve is almost exactly consistent with the theoretical range.
-Furthermore, assume that $\{R_{\text{CoT}}^{i}\}_{i=1}^{\infty}$ is bounded thus we can easily prove that the overall resistance has a lower bound:
-$\lim_{n\rightarrow+\infty}R^{\prime}_{\text{all}}=R_{0}+R_{S}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}=R_{0}+R_{S}.$ (15)
+$P^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}})^{2}}.$ (16)
+As the number of samples increases, the curves in Fig.
+5c reveal that the real performance curve is almost exactly consistent with the theoretical range.
+Furthermore, assume that $\{R_{\text{ITR}}^{i}\}_{i=1}^{\infty}$ is bounded thus we can easily prove that the overall resistance has a lower bound:
+$\lim_{n\rightarrow+\infty}R^{\prime}_{\text{all}}=R_{0}+R_{S}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}=R_{0}+R_{S}.$ (17)
 The corresponding performance bound is:
-$\hat{P}^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{0}+R_{S})^{2}}$ (16)
-As shown in Figure 4c, model performance aligns with this theoretical upper bound as sample size increases.
+$\hat{P}^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{0}+R_{S})^{2}}$ (18)
+As shown in Fig. 5c, model performance aligns with this theoretical upper bound as sample size increases.
 B.1.2 Analysis of Effectiveness of Coverage Design
@@
 Under this assumption, the integration difficulty in self-consistency verification approaches zero.
-Analogously, this can be modeled as a zero-value combined resistance in parallel with multiple CoT resistors, where the coverage resistance of self-consistency $\hat{R}_{\text{all}}$ is given by:
-$\hat{R}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}=R_{0}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}$ (17)
+Analogously, this can be modeled as a zero-value combined resistance in parallel with multiple ITR resistors, where the coverage resistance of self-consistency $\hat{R}_{\text{all}}$ is given by:
+$\hat{R}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}=R_{0}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}$ (19)
 where $n$ denotes the number of samples. Using this, the theoretical output power can be similarly calculated as:
-$P^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(\frac{1}{\sum_{n}\frac{1}{R_{\text{CoT}}}}+R_{0})^{2}},$ (18)
+$P^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(\frac{1}{\sum_{n}\frac{1}{R_{\text{ITR}}}}+R_{0})^{2}},$ (20)
 which are much more smaller than $P^{\text{self}}_{\text{out}}$ in the same $n$ sample size with $R_{S}\gg 0$.
-As the number of samples increases, the curve in Figure 4c demonstrates that the real performance aligns closely with the theoretical range.
-Building upon the same bounded assumption of $\{R_{\text{CoT}}^{i}\}_{i=1}^{\infty}$, this setup also defines a resistance lower bound:
-$\lim_{n\rightarrow+\infty}\hat{R}_{\text{all}}=R_{0}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}=R_{0}$ (19)
+As the number of samples increases, the curve in Fig.
+5c demonstrates that the real performance aligns closely with the theoretical range.
+Building upon the same bounded assumption of $\{R_{\text{ITR}}^{i}\}_{i=1}^{\infty}$, this setup also defines a resistance lower bound:
+$\lim_{n\rightarrow+\infty}\hat{R}_{\text{all}}=R_{0}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}=R_{0}$ (21)
 and the corresponding lower-bound performance:
-$\hat{P}^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}}{R_{0}}.$ (20)
-As shown in Figure 4c, the model’s performance gradually converges toward this upper bound and much larger than the performance of self-consistency, demonstrating consistency with the theoretical predictions.
+$\hat{P}^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}}{R_{0}}.$ (22)
+As shown in Fig.
+5c, the model’s performance gradually converges toward this upper bound and much larger than the performance of self-consistency, demonstrating consistency with the theoretical predictions.
 NOTE
 According to the theoretical formulation in Eq.
-18, in the reasoning phase where the resistance is significantly greater than $R_{0}$, the overall trend shows that Coverage is approximately proportional to the number of samples.
+20, in the reasoning phase where the resistance is significantly greater than $R_{0}$, the overall trend shows that Coverage is approximately proportional to the number of samples.
 Meanwhile, certain Inference Scaling Laws indicate that Coverage and the logarithm of the number of samples exhibit an approximately proportional relationship.
-These two observations are not contradictory, as our model assumes relatively independent sampling.
+These two observations are not contradictory, as ECP assumes relatively independent sampling.
 However, we have observed that when the number of samples exceeds 100, an increasing proportion of the samples become duplicates rather than independent samples.
@@
 According to Eq. 15, the original circuit can be considered as a large parallel reasoning and verification process and the total resistor is as follows:
-$R^{\prime}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}.$ (21)
-As shown in Figure 6a, this process is decomposed into multiple sequential, smaller-scale parallel reasoning and verifications.
+$R^{\prime}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}.$ (23)
+As shown in Fig.
+7a, this process is decomposed into multiple sequential, smaller-scale parallel reasoning and verifications.
 This refinement transforms a single large parallel reasoning resistance with significant verification resistance into a sequence of smaller parallel reasoning resistance and smaller verification resistance.
@@
 where $n$ denotes the sampling size, $R_{j}^{S}$ represents the small and fine-grained verification resistance for $j$-th step, and $R_{j}^{i}$ corresponds to smaller reasoning resistance for $j$-th step for $i$-th sampling.
-The total resistance can be broken down into resistance of sub-step resistor as $R^{i}_{\text{CoT}}=\sum_{j}R_{j}^{i}$.
+The total resistance can be broken down into resistance of sub-step resistor as $R^{i}_{\text{ITR}}=\sum_{j}R_{j}^{i}$.
 We aim to demonstrate that our approach indeed optimizes the total resistance, i.e., $R^{\text{fine}}_{\text{all}}<R^{\prime}_{\text{all}}$.
@@
 The proof is as follows:
-$R^{\text{fine}}_{\text{all}}-R^{\prime}_{\text{all}}$ $=\sum_{j}R_{j}^{S}-R_{S}+\frac{R_{\text{CoT}}}{n}-\sum_{j}\frac{R_{j}^{i}}{n}$ (23)
+$R^{\text{fine}}_{\text{all}}-R^{\prime}_{\text{all}}$ $=\sum_{j}R_{j}^{S}-R_{S}+\frac{R_{\text{ITR}}}{n}-\sum_{j}\frac{R_{j}^{i}}{n}$ (25)
 $=\sum_{j}R_{j}^{S}-R_{S}.$ (26)
@@
 The original equation is thus reformulated as:
-$R^{\text{CoV}}_{\text{all}}$ $=R_{0}+\frac{1}{\sum_{j=1}^{k}\sum_{i=1}^{n}\frac{1}{R_{\text{S}}}}+\sum_{j=1}^{k}\left(R_{\text{meta}}\right)+\frac{R_{\text{CoT}}}{n}$
-$=R_{0}+\frac{R_{\text{S}}}{nk}+kR_{\text{meta}}+\frac{R_{\text{CoT}}}{n}.$
+$R^{\text{CoV}}_{\text{all}}$ $=R_{0}+\frac{1}{\sum_{j=1}^{k}\sum_{i=1}^{n}\frac{1}{R_{\text{S}}}}+\sum_{j=1}^{k}\left(R_{\text{meta}}\right)+\frac{R_{\text{ITR}}}{n}$
+$=R_{0}+\frac{R_{\text{S}}}{nk}+kR_{\text{meta}}+\frac{R_{\text{ITR}}}{n}.$
 Clearly, as the sample size $n$ and the meta-validation steps $k$ increase, $\frac{R_{\text{S}}}{nk}$ decreases, leading to reduced difficulty in meta-validation.
 This significantly optimizes the total resistance and improves output power.
-Figure 6b illustrates this performance enhancement.
+Fig.
+7b illustrates this performance enhancement.
 However, as the number of samples requiring validation grows, the model’s accuracy first increases and then decreases. Specifically, for sufficiently large validation steps, the resistance satisfies:
@@
 Consequently, $R^{\text{CoV}}_{\text{all}}$ exceeds its optimized value, leading to a performance trend characterized by initial improvement followed by decline.
-C Proof of Impact of Resistor Optimization
-To compute the derivatives of $P_{\text{model}}$ with respect to $R_{\text{CoT}}$, we start with the given expression:
-$P_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{\text{CoT}}+R_{0})^{2}}.$ (26)
-According to Eq. 10, let $R_{\text{CoT}}=R_{1}+R_{2}$, where $R_{1}\gg R_{2}$, if $R_{1}$ is improved to $\frac{R_{1}}{k}$, the power can be calculated as:
-$P^{1}_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (27)
+B.4 Proof of Impact of Resistor Optimization
+To compute the derivatives of $P_{\text{model}}$ with respect to $R_{\text{ITR}}$, we start with the given expression:
+$P_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{\text{ITR}}+R_{0})^{2}}.$ (28)
+According to Eq. 11, let $R_{\text{ITR}}=R_{1}+R_{2}$, where $R_{1}\gg R_{2}$, if $R_{1}$ is improved to $\frac{R_{1}}{k}$, the power can be calculated as:
+$P^{1}_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (29)
 The corresponding power improvement is:
 $\Delta P^{1}_{\text{model}}$ $=P^{1}_{\text{model}}-P_{\text{model}}$ (30)
-$=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (29)
-$=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}})$ (30)
+$=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (31)
+$=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}})$ (32)
 Similarly, if $R_{2}$ is improved to the original $\frac{R_{2}}{k}$, the power improvement can be calculated as:
-$\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (31)
+$\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (33)
 To determine which improvement has a greater effect, consider:
-$\Delta P^{1}_{\text{model}}-\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}(\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (32)
+$\Delta P^{1}_{\text{model}}-\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}(\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (34)
 Define:
@@
 $B^{2}-A^{2}=(B-A)(B+A)$ (43)
-We only need to examine the sign of $B-A$, since $B+A>0$ holds true at all times.since $R_{1}\gg R_{2}$, and $01-\frac{1}{k}>$ 0 for $n>1$, we have:
+We only need to examine the sign of BAB-A, since $B+A>0$ holds true at all times.since $R_{1}\gg R_{2}$, and $1-\frac{1}{k}>0$ for $n>1$, we have:
 $R_{1}\left(1-\frac{1}{k}\right)-R_{2}\left(1-\frac{1}{k}\right)$ $>0$ (44)